### v2

# A torsion theoretic interpretation for sheaves of modules and Grothendieck topologies on directed categories

Abstract.
We prove that every Grothendieck topology induces a hereditary torsion pair in the category of presheaves of modules on a ringed site, and obtain a homological characterization of sheaves of modules: a presheaf of modules is a sheaf of modules if and only if it is saturated with respect to torsion presheaves, or equivalently, it is right perpendicular to torsion presheaves in the sense of Geigle and Lenzing. We also study Grothendieck topologies on directed categories ${\mathscr{C}}$, and show that every Grothendieck topology on ${\mathscr{C}}$ is a subcategory topology if and only if ${\mathscr{C}}$ is an artinian EI category. Consequently, in this case every sheaf category is equivalent to the presheaf category over a full subcategory of ${\mathscr{C}}$. Finally, we classify all Grothendieck topologies on a special type of noetherian EI categories, and extend the locally self-injective property of representations of ${\mathrm{FI}}$ and ${\mathrm{VI}}$ to representations of their infinite full subcategories. Some potential applications in group representation theory are given at the end of this paper.
Key words and phrases:
Grothendieck topologies, sheaves, torsion pairs, sheafification, sheaf cohomology, Serre quotient, directed categories, EI categories
Z.X. Di was partly supported by NSF of China (Grant No. 11971388), the Scientific Research Funds of Huaqiao University (Grant No. 605-50Y22050) and the Fujian Alliance of Mathematics (Grant No. 2024SXLMMS04); L.P. Li was partly supported by NSF of China (Grant No. 12171146); L. Liang was partly supported by NSF of China (Grant No. 12271230).
1. Introduction
Sheaves of modules on ringed sites have been widely applied in algebraic geometry, algebraic topology, and geometric representation theory. Recently, they were also used to investigate modular representation theory of finite groups [3, 25, 26, 27] and continuous representations of topological groups [7]. This new approach combines ideas and methods from topos theory, representation theory of categories, and representation theory of groups, and hence not only proposes a comprehensive technique to study representations of groups, but also reveals deep relations among distinct areas. A classical example is Artin’s theorem [18, III. 9, Theorems 1 and 2], from which it follows that the category of continuous representations of a topological group $G$ is equivalent to the category of sheaves of modules over a ringed site whose underlying category is an orbit category of $G$ equipped with the atomic Grothendieck topology.
Sheaf theory is expected to play a more significant role in algebraic representation theory. To facilitate its applications in this area, it is useful to reinterpret concepts and notions in topos theory, which are formal and sometimes hard to check in practice, in terms of representation theory. In [7] we give an attempt for the special case of the atomic Grothendieck topology via a particular torsion theory. The first goal of this paper is to extend it to the full generality of all Grothendieck topologies.
Let us give some details. Let ${\mathscr{C}}$ be a small category, $J$ a Grothendieck topology on the opposite category ${\mathscr{C}^{\mathrm{op}}}$, [^1] In this paper we consider representations of ${\mathscr{C}}$; that is, covariant functors from ${\mathscr{C}}$ to a concrete category. Since they are precisely presheaves over ${\mathscr{C}^{\mathrm{op}}}$, we impose Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ rather than on ${\mathscr{C}}$. and $\mathcal{O}$ a structure sheaf over the site $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$. Given a presheaf $V$ of modules over the ringed site $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ (also called an $\mathcal{O}$-module) as well as an object $x$ in ${\mathscr{C}}$, we define an element $v\in V_{x}$, the value of $V$ on $x$, to be $J$-torsion if there is a covering sieve $S\in J(x)$ such that every morphism in $S$ sends $v$ to 0. The axioms of Grothendieck topologies guarantee that all $J$-torsion elements in $V$ form an $\mathcal{O}$-submodule; see Lemma 3.2. Consequently, the operation to take the maximal $J$-torsion submodule gives rise to a left exact endo-functor $\mathcal{T}_{J}$ on $\mathcal{O}\operatorname{-Mod}$, the category of $\mathcal{O}$-modules. We call this functor the torsion functor and say that $V$ is $J$-torsion (resp., $J$-torsion free) if $\mathcal{T}_{J}(V)=V$ (resp., $\mathcal{T}_{J}(V)=0$). If $V$ is torsion free and $\mathrm{R}^{1}\mathcal{T}_{J}(V)=0$, where $\mathrm{R}^{1}\mathcal{T}_{J}$ is the first right derived functor of $\mathcal{T}_{J}$, then we call it a $J$-saturated module.
Denote by $\mathcal{T}(J)$ (resp., $\mathcal{F}(J)$) the category of $J$-torsion (resp., $J$-torsion free) $\mathcal{O}$-modules. We show that $(\mathcal{T}(J),\mathcal{F}(J))$ is a hereditary torsion pair on $\mathcal{O}\operatorname{-Mod}$, and it completely determines the Grothendieck topology. That is, if $K$ is another Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ with respect to which $\mathcal{O}$ is still a structure sheaf, then $J=K$ if and only if $\mathcal{T}(J)=\mathcal{T}(K)$. For details, see Proposition 3.8 and Proposition 3.13.
The following result gives a characterization of sheaves of modules over $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ in terms of the torsion theory mentioned above, and extends [7, Theorem 3.8] to full generality. Compared to the original definition of sheaves in terms of amalgamations of matching families, this characterization is homological, and is more elementary for applications in representation theory.
Theorem 1.1.
An $\mathcal{O}$-module $V$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ if and only if it is $J$-saturated.
It is direct to check that $J$-saturated $\mathcal{O}$-modules are precisely $\mathcal{O}$-modules right perpendicular to $J$-torsion $\mathcal{O}$-modules in the sense of Geigle and Lenzing [12]; for details, see [12, Section 1]. Consequently, applying [12, Propositions 2.2 and 2.5], one immediately deduces the following result:
Corollary 1.2.
One has the following equivalence
${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})\simeq\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J),$
namely the category of sheaves of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is equivalent to the Serre quotient of the category of $\mathcal{O}$-modules by the category of $J$-torsion $\mathcal{O}$-modules.
Remark 1.3.
It is well known that the sheaf category is a localization of the presheaf category by a localizing subcategory. This corollary gives an explicit description of the localizing subcategory: the full subcategory consists of all $J$-torsion modules.
The above homological characterization of sheaves of modules and equivalence of categories bring quite a few useful consequences. For instance, injective objects in the sheaf category are precisely $J$-torsion free injective $\mathcal{O}$-modules (see Corollary 4.4); sheaf cohomology can be reinterpreted via the torsion functor $\mathcal{T}_{J}$ (see Corollary 4.6). Furthermore, one can obtain a very elementary description of the sheafification functor (see the explanation after Corollary 4.5).
The Serre quotient still seems mysterious for practical purpose. For instance, even for the special case that ${\mathscr{C}^{\mathrm{op}}}$ is an orbit category of a finite group, $J$ is the atomic Grothendieck topology, and $\mathcal{O}$ is a constant structure sheaf induced by a field, in general it is a hard task to classify irreducible objects in $\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J)$ (as we will see later, an answer of this question can bring a significant progress for verifying Alperin’s weight conjecture). The second goal of this paper is to find some specific categories such that one can obtain a relatively explicit description of this Serre quotient.
It has been pointed out in [2, IV.9] by Grothendieck and Verdier that every Grothendick topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid (or called a subcategory topology) when ${\mathscr{C}}$ is a finite category, and in this case ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is equivalent to the category of presheaves of modules on a full subcategory of ${\mathscr{C}^{\mathrm{op}}}$. Similar results have been established by Hemelaer in [14] and Lindenhovius in [17] for posets satisfying certain extra conditions. Although it seems to be a very difficult task to classify small categories whose Grothendieck topologies are rigid, in this paper we obtain a complete answer for categories satisfying the following combinatorial condition: the binary relation $\leqslant$ on $\operatorname{Ob}({\mathscr{C}})$ defined by setting $x\leqslant y$ if and only if ${\mathscr{C}}(x,y)\neq\emptyset$ is a partial order. We call them directed categories. The reason we choose them is plain: many important categories of particular interest in representation theory are directed categories, for examples, posets, orbit categories of groups, quivers without loops or oriented cycles, and etc.
Theorem 1.4.
Let ${\mathscr{C}}$ be a directed category. Then every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if ${\mathscr{C}}$ is a noetherian EI category; that is, the poset $(\operatorname{Ob}({\mathscr{C}}),\leqslant)$ satisfies the ascending chain condition and every endomorphism in ${\mathscr{C}}$ is an isomorphism.
Consequently, when ${\mathscr{C}}$ is a noetherian EI category, questions concerning ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ can be completely transferred to questions about representations of full subcategories of ${\mathscr{C}}$, so we have no “generic” sheaf theory in this situation. Explicitly, by the Comparison Lemma [13, C2.2, Theorem 2.2.3] (see also [13, C2.2, Example 2.2.4(d)]), one has:
Corollary 1.5.
Let ${\mathscr{C}}$ be a noetherian EI category, $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$, $\mathcal{O}$ a structure sheaf over $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$, and $\mathscr{D}$ the full subcategory of ${\mathscr{C}}$ consisting of objects $x$ with $J(x)=\{{\mathscr{C}}(x,-)\}$. Then:
  1. (1)
    an $\mathcal{O}$-module $V$ lies in $\mathcal{T}(J)$ if and only if $V_{x}=0$ for all $x\in\operatorname{Ob}({\mathscr{D}})$;
  2. (2)
    one has
    ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})\simeq\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod},$
    where $\mathcal{O}_{{\mathscr{D}}}$ is the structure sheaf on ${\mathscr{D}}^{{\mathrm{op}}}$ obtained via restricting $\mathcal{O}$ to ${\mathscr{D}}^{{\mathrm{op}}}$.
However, when ${\mathscr{C}}$ is not noetherian, a classification of Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ becomes much harder. In this paper we solve this question only for EI categories ${\mathscr{C}}$ of type $\mathbb{N}$ (see the beginning of Section 6 for a precise definition) or type $\mathbb{Z}$. Although it seems to the reader that too many combinatorial restrictions are imposed on them, there are still quite a few interesting examples whose representation theory are actively studied (see for instances [6, 20, 21]):
  • skeletal full subcategories of the category ${\mathrm{FI}}$ of finite sets and injections, where ${\mathrm{FI}}$ is equivalent to the opposite category of an orbit category of the permutation group on ${\mathbb{N}}$ ([13, Examples A2.1.11(h) and D3.4.10]);
  • skeletal full subcategories of the opposite category of the category $\mathrm{FS}$ of finite sets and surjections, where $\mathrm{FS}$ is equivalent to an orbit category of the automorphism group of the free Boolean algebra on a countable infinity of generators, or the self-homeomorphism group of the Cantor set ([13, Example D3.4.12]);
[…]
We can classify all Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ when ${\mathscr{C}}$ is an EI category of type $\mathbb{N}$ or $\mathbb{Z}$.
Theorem 1.6.
Let ${\mathscr{C}}$ be an EI category of type $\mathbb{N}$ or $\mathbb{Z}$. One has:
  1. (1)
    Grothendieck topologies $J$ on ${\mathscr{C}^{\mathrm{op}}}$ such that $J(x)$ does not contain the empty functor for each $x\in\operatorname{Ob}({\mathscr{C}})$ are parameterized by functions
    $\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})\longrightarrow{\mathbb{N}}\cup\{\infty\}$
    satisfying the following condition: if $\boldsymbol{d}(n)\neq 0$, then $\boldsymbol{d}(n+1)=\boldsymbol{d}(n)-1$.
  2. (2)
    All other Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by functions
    $\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})\longrightarrow{\mathbb{N}}\cup\{-\infty\}$
    satisfying the following conditions: $\boldsymbol{d}(n)=-\infty$ for $n\gg 0$; and if $\boldsymbol{d}(n)\neq 0$, then $\boldsymbol{d}(n+1)=\boldsymbol{d}(n)-1$.
  3. (3)
    A Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ is rigid if and only if its corresponded function $\boldsymbol{d}$ satisfies the following condition: $\boldsymbol{d}(n)\neq\infty$ for all $n\in{\mathbb{N}}$.
For an EI category ${\mathscr{C}}$ of type $\mathbb{N}$ or $\mathbb{Z}$, it follows that a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ which is not a subcategory topology is almost atomic; that is, there is a full subcategory of ${\mathscr{C}^{\mathrm{op}}}$ containing all but finitely many objects such that the restriction of $J$ to this subcategory is the atomic Grothendieck topology. Accordingly, to describe sheaves of modules over arbitrary ringed sites $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})$, we only need to understand sheaves of modules over ringed atomic sites. For a precise meaning of this statement, see Proposition 6.9 and Remark 6.10.
We can use Theorem 1.6 and the Comparison Lemma to deduce some interesting results in representation theory of combinatorial categories. For instance, let $k$ be a field of characteristic 0, and let ${\mathscr{C}}$ be the category of finite sets and injections, or the category of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved in [11] (see also [22]) that every indecomposable projective representation of ${\mathscr{C}}$ over $k$ is also injective. It turns out this locally self-injective property holds for all (essentially) infinite full subcategories of ${\mathscr{C}}$.
Theorem 1.7.
Let ${\mathscr{C}}$ be specified as above, $k$ a field of characteristic 0, and ${\mathscr{D}}$ a full subcategory of ${\mathscr{C}}$ containing infinitely many objects (up to isomorphism). Then every indecomposable projective ${\mathscr{D}}$-module over $k$ is also injective.
For a generalized version of this result, see Theorem 7.1.
This paper is organized as follows. In Section 2 for the reader’s convenience we provide some background knowledge on Grothendieck topologies and sheaf theory. In Section 3 we construct a torsion theory induced by a rule $J$ satisfying partial properties of Grothendieck topology. Indeed, for most results described in this section we do not require $J$ to be a Grothendieck topology, so the reader can understand the role of each individual property. The main content of Section 4 consists of a proof of Theorem 1.1 and several consequences. Directed categories are introduced in Section 5, where we also give a proof of Theorem 1.4. In Section 6 we classify Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ when ${\mathscr{C}}$ is an EI category of type ${\mathbb{N}}$, proving Theorem 1.6. In the last section we describe some potential applications of previously established results and prove Theorem 1.7.
Throughout this paper, unless otherwise specified, all functors are covariant, modules over rings are left modules, and composition of maps and morphisms is from right to left.
2. Preliminaries
Let ${\mathscr{C}}$ be a small skeletal category and let ${\mathscr{C}^{\mathrm{op}}}$ be its opposite category. In this section we briefly recall definitions and elementary facts on Grothendieck topologies and sheaves. For more details, please refer to [1, 2, 13, 16, 18, 23].
2.1. Grothendieck topologies
Let $x$ be an object in ${\mathscr{C}}$. A left ideal $S$ of morphisms on $x$ is a subfunctor of the representable functor ${\mathscr{C}}(x,-)$. Equivalently, $S$ is a set of morphisms with domain $x$ such that the following condition holds: if $f:x\to y$ is a member in $S$, and $g:y\to z$ is an arbitrary morphism in ${\mathscr{C}}$, then $gf=g\circ f$ is contained in $S$ as well. A left ideal of morphisms on $x$ is also called a sieve on $x$ in ${\mathscr{C}^{\mathrm{op}}}$.
Throughout this paper we fix $J$ to be a rule assigning to each object $x$ a collection $J(x)$ of left ideals of morphisms on $x$. The following definition is taken from [18, III.2, Definition 1], which defines a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ if $J$ satisfies all three axioms. Here we use the dual version, since we mainly consider representations of ${\mathscr{C}}$ which are covariant functors, rather than presheaves which are contravariant functors.
Definition 2.1.
We say that $J$ satisfies
  1. (1)
    the maximal axiom if ${\mathscr{C}}(x,-)\in J(x)$ for each $x\in\operatorname{Ob}({\mathscr{C}})$;
  2. (2)
    the stability axiom if $S\in J(x)$ implies that the pullback
    $f^{\ast}(S)=\{g:y\to\bullet\mid gf\in S\}\in J(y)$
    for any morphism $f:x\to y$ in ${\mathscr{C}}$;
  3. (3)
    the transitivity axiom if $S\in J(x)$ and $T$ is a left ideal of morphisms on $x$ such that $f^{\ast}(T)\in J(y)$ for any $(f:x\to y)\in S$, then $T\in J(x)$.
The following fact is well known; see [18, Section III.2]. For the convenience of the reader, we give a complete proof.
Lemma 2.2.
If $J$ is a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$, then:
  1. (1)
    $J$ is closed under inclusions: if $S_{1}$ is a subfunctor of ${\mathscr{C}}(x,-)$, $S_{2}$ is a subfunctor of $S_{1}$ and $S_{2}\in J(x)$, then $S_{1}\in J(x)$ as well;
  2. (2)
    $J$ is closed under finite intersections: if $S_{1},S_{2}\in J(x)$, then $S_{1}\cap S_{2}\in J(x)$.
Proof.
For every $f:x\to y$ in $S_{2}$, one has $f^{\ast}(S_{1})={\mathscr{C}}(y,-)\in J(y)$, so (1) follows from the transitivity axiom. Similarly, for every $f:x\to y$ in $S_{2}$, one can check that $f^{\ast}(S_{1}\cap S_{2})=f^{\ast}(S_{1})$, so $f^{\ast}(S_{1}\cap S_{2})$ is contained in $J(y)$ by the stability axiom. Consequently, $S_{1}\cap S_{2}\in J(x)$ by the transitivity axiom. ∎
Example 2.3.
We list a few important Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ below:
  • the trivial topology $J_{t}$: $J_{t}(x)=\{{\mathscr{C}}(x,-)\}$ for every $x\in\operatorname{Ob}({\mathscr{C}})$;
  • the maximal topology $J_{m}$: $J_{m}(x)$ consists of all subfunctors (including $\emptyset$) of ${\mathscr{C}}(x,-)$;
  • the dense topology $J_{d}$: a subfunctor $S\subseteq{\mathscr{C}}(x,-)$ is contained in $J_{d}(x)$ if for every morphism $f:x\to y$, there exists a morphism $g:y\to z$ such that $gf\in S$;
  • the atomic Grothendieck topology $J_{a}$: $J_{a}(x)$ consists of all nonempty subfunctors of ${\mathscr{C}}(x,-)$ provided that ${\mathscr{C}}$ satisfies the following Ore condition: given two morphisms $f:x\to y$ and $g:x\to z$, there exist morphisms $f^{\prime}:y\to w$ and $g^{\prime}:z\to w$ such that $f^{\prime}f=g^{\prime}g$.
Note that the empty subfunctor $\emptyset$ can never belong to $J_{d}(x)$ for $x\in\operatorname{Ob}({\mathscr{C}})$. Furthermore, if ${\mathscr{C}}$ satisfies the right Ore condition, then $J_{d}=J_{a}$.
2.2. Sheaf theory
Fix a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ and let $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ be the corresponded site. The following definitions are taken from [18, III.4].
Definition 2.4.
Let ${\mathrm{Set}}$ be the category of sets and maps.
  • A presheaf of sets on ${\mathscr{C}^{\mathrm{op}}}$ is a covariant functor $F:{\mathscr{C}}\to{\mathrm{Set}}$.
  • Given $S\in J(x)$, a matching family for $S$ of elements of $F$ is a rule assigning to each morphism $f:x\to y$ in $S$ an element $v_{f}\in F(y)$ such that $g\cdot v_{f}=v_{gf}$ for any morphism $g:y\to\bullet$, where the action of $g$ on $v_{f}$ is determined by the covariant functor $F$.
  • An amalgamation of the above matching family is an element $v_{x}\in F(x)$ such that $f\cdot v_{x}=v_{f}$ for all $f\in S$.
More formally, let $\mathscr{S}$ be the category whose objects are elements in $S$ and morphisms are dotted arrows making the following diagram commutes
${x\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\in S}$$\scriptstyle{g\in S}$${y}$${z\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$
Let $\mathscr{S}_{\ast}$ be the larger category obtained by adding the identity morphism $1_{x}$ on $x$ to $\mathscr{S}$. One can check that a matching family is a functor $M$ from $\mathscr{S}$ to the category of singleton sets such that $M(f)\subseteq(F\circ\operatorname{cod})(f)$, where $\operatorname{cod}$ is the codomain functor, and an amalgamation of this matching family is an extension of $M$ from $\mathscr{S}$ to $\mathscr{S}_{\ast}$.
Definition 2.5.
A presheaf $F$ of sets on ${\mathscr{C}^{\mathrm{op}}}$ is called separated if every matching family for every object $x$ and every $S\in J(x)$ has at most one amalgamation. It is a sheaf of sets on the site $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ if every matching family has exactly one amalgamation.
Remark 2.6.
A more concise version of this definition is: $F$ is a sheaf of sets on the site $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$ if and only if for every $x\in\operatorname{Ob}({\mathscr{C}})$ and every $S\in J(x)$, the inclusion $S\subseteq{\mathscr{C}}(x,-)$ induces an isomorphism
$\mathrm{Nat}({\mathscr{C}}(x,-),F)\cong\mathrm{Nat}(S,F)$
where $\mathrm{Nat}(\bullet,F)$ means the set of natural transformations from $\bullet$ to $F$.
When $J$ is the trivial topology, every presheaf of sets on ${\mathscr{C}^{\mathrm{op}}}$ is a sheaf of sets. For a general Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$, the category ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$ of sheaves of sets on $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$ is a full subcategory of ${\mathscr{C}}\operatorname{-Set}$, the category of all presheaves of sets on ${\mathscr{C}^{\mathrm{op}}}$. The inclusion functor $\iota:{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})\to{\mathscr{C}}\operatorname{-Set}$ has a left adjoint $\sharp:{\mathscr{C}}\operatorname{-Set}\to{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$, called the sheafification functor. It is well known that the sheafification is a special type of localization; see [10, Proposition 1.3] or [15]. Therefore, ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}})$ is a localization of ${\mathscr{C}}\operatorname{-Set}$ by the multiplicative system of local isomorphisms [15] (or morphismes bicouvrants [2]).
Now we give definitions of ringed sites and sheaves of modules on ringed sites. For more details, please refer to [23, Chapter 18].
Definition 2.7.
Let $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}=({\mathscr{C}^{\mathrm{op}}},J)$ be a site.
  1. (1)
    A presheaf of rings on $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$ is a covariant functor $\mathcal{O}:{\mathscr{C}}\to\mathrm{Ring}$, the category of associative rings, and it is called a sheaf of rings if the underlying presheaf of sets is a sheaf of sets.
  2. (2)
    If $\mathcal{O}$ is a sheaf of rings on $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$, then the pair $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is called a ringed site, and $\mathcal{O}$ is called the structure sheaf.
  3. (3)
    A presheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is covariant functor $V:{\mathscr{C}}\to\mathrm{Ab}$, the category of abelian groups, satisfying the following conditions: the value $V_{x}$ of $V$ on $x$ is an $\mathcal{O}_{x}$-module for each $x\in\operatorname{Ob}({\mathscr{C}})$, and $V_{f}:V_{x}\to V_{y}$ is an $\mathcal{O}_{x}$-linear map for each morphism $f:x\to y$ in ${\mathscr{C}}$, where $V_{y}$ is viewed as an $\mathcal{O}_{x}$-module by the ring homomorphism $\mathcal{O}_{f}:\mathcal{O}_{x}\to\mathcal{O}_{y}$. A presheaf of modules is called a sheaf of modules if the underlying presheaf of sets is a sheaf of sets.
In this paper we also call a presheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ an $\mathcal{O}$-module. Denote the category of all $\mathcal{O}$-modules by $\mathcal{O}\operatorname{-Mod}$ and the category of sheaves of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ by ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$. Note that ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is an abelian category: kernels are computed in $\mathcal{O}\operatorname{-Mod}$; to compute cokernels, we first obtain cokernels in $\mathcal{O}\operatorname{-Mod}$, and then apply the sheafification functor $\sharp$. Furthermore, ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is a Grothendieck category, and hence every object in it has an injective hull. But in general it does not have enough projective objects.
3. $J$-torsion theory
In this section we develop a torsion theory induced by rules $J$ satisfying certain conditions. As before, let ${\mathscr{C}}$ be a small skeletal category, and let $\mathcal{O}$ be a structure presheaf, namely a presheaf of associative rings on ${\mathscr{C}^{\mathrm{op}}}$. [^2] The reason we consider structure presheaves rather than structure sheaves is that $J$ might not be a Grothendieck topology, and in that case structure sheaves do not make sense. We study $J$-torsion $\mathcal{O}$-modules, and describe a bijective correspondence between Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ and certain special hereditary torsion pairs on $\mathcal{O}\operatorname{-Mod}$. We also note that the results in this section parallel the torsion theory of module categories induced by Gabriel topologies on associative rings; for details, please see [4] or [5, Section 4].
3.1. $J$-torsion modules
The main content of this subsection includes elementary properties of $J$-torsion $\mathcal{O}$-modules. Let $V$ be an $\mathcal{O}$-module. We define $J$-torsion elements of $V$ as follows.
Definition 3.1.
For $x\in\operatorname{Ob}({\mathscr{C}})$, an element $v\in V_{x}$ is called $J$-torsion if there is a certain $S\in J(x)$ such that $S\cdot v=0$; that is, $f\cdot v=V_{f}(v)=0$ for each $f\in S$.
Given $S\in J(x)$, we set
$\mathcal{T}_{S}(V)=\{v\in V_{x}\mid S\cdot v=0\},$
which is an $\mathcal{O}_{x}$-submodule of $V_{x}$. In particular, if $\emptyset\in J(x)$, then $\mathcal{T}_{\emptyset}(V)=V_{x}$ by convention, so every $v\in V_{x}$ is trivially $J$-torsion in this case.
Now we define
$\mathcal{T}_{J}(V)=\bigoplus_{x\in\operatorname{Ob}({\mathscr{C}})}\sum_{S\in J(x)}\mathcal{T}_{S}(V).$
At this moment it is not clear that $\mathcal{T}_{J}(V)$ is an $\mathcal{O}$-submodule of $V$. However, when $J$ satisfies the stability axiom, it is indeed the case.
Lemma 3.2.
Notation as above and suppose that $J$ satisfies the stability axiom. Then:
  1. (1)
    $\mathcal{T}_{J}(V)$ is an $\mathcal{O}$-submodule of $V$;
  2. (2)
    the rule $V\to\mathcal{T}_{J}(V)$ is a left exact functor from $\mathcal{O}\operatorname{-Mod}$ to itself;
  3. (3)
    if $J$ is closed under finite intersections, then every element in $(\mathcal{T}_{J}(V))_{x}$ is $J$-torsion for $x\in\operatorname{Ob}({\mathscr{C}})$.
Proof.
(1): It suffices to check the following condition: for each morphism $f:x\to y$ in ${\mathscr{C}}$,
$v\in\sum_{S\in J(x)}\mathcal{T}_{S}(V)\Rightarrow f\cdot v\in\sum_{S\in J(y)}\mathcal{T}_{S}(V).$
Write $v=v_{1}+v_{2}+\ldots+v_{s}$ such that each $v_{i}$ is a $J$-torsion element, meaning that there is a certain $S_{i}\in J(x)$ (the special case $S_{i}=\emptyset$ is allowed) satisfying $S_{i}\cdot v_{i}=0$ for $1\leqslant i\leqslant s$. Clearly, $f^{\ast}(S_{i})\cdot(f\cdot v_{i})=0$. But $f^{\ast}(S_{i})\in J(y)$ by the stability axiom, so
$f\cdot v_{i}\in\mathcal{T}_{f^{\ast}(S_{i})}(V)\subseteq\sum_{S\in J(y)}\mathcal{T}_{S}(V)$
for each $i$. Consequently, $f\cdot v$ is contained in $\sum_{S\in J(y)}\mathcal{T}_{S}(V)$ as desired.
(2): We want to show that $\mathcal{T}_{J}$ sends a morphism $\phi:V\to W$ between two $\mathcal{O}$-modules to a morphism $\mathcal{T}_{J}(V)\to\mathcal{T}_{J}(W)$ via restriction. It suffices to explain that $\phi_{x}$ sends $J$-torsion elements in $V_{x}$ to $J$-torsion elements in $W_{x}$ for $x\in\operatorname{Ob}({\mathscr{C}})$. Let $v\in V_{x}$ be a $J$-torsion element. If $\emptyset\in J(x)$, then every element in $W_{x}$ is $J$-torsion, and the conclusion holds trivially. Otherwise, there is a nonempty $S\in J(x)$ such that $f\cdot v=0$ for every $(f:x\to y)\in S$. Consequently, $f\cdot\phi_{x}(v)=\phi_{y}(f\cdot v)=0$, so $\phi_{x}(v)$ is $J$-torsion as well.
(3): Take an arbitrary $v\in(\mathcal{T}_{J}(V))_{x}$ and write it as $v_{1}+\ldots+v_{s}$ such that $v_{i}\in\mathcal{T}_{S_{i}}(V)$ for a certain $S_{i}\in J(x)$. Let $S=S_{1}\cap\ldots\cap S_{s}$ which is contained in $J(x)$ since $J$ is closed under finite intersections. Then $v\in\mathcal{T}_{S}(V)$, and hence is a $J$-torsion element. ∎
Setup: In the rest of this paper we always assume that $J$ satisfies the stability axiom.
Definition 3.3.
We call $\mathcal{T}_{J}(V)$ the $J$-torsion submodule of $V$. We say that $V$ is
  • $J$-torsion if $\mathcal{T}_{J}(V)=V$;
  • $J$-torsion free if $\mathcal{T}_{J}(V)=0$;
  • $J$-saturated if $\mathrm{R}^{i}\mathcal{T}_{J}(V)=0$ for $i\leqslant 1$, where $\mathrm{R}^{i}\mathcal{T}_{J}$ is the $i$-th right derived functor of $\mathcal{T}_{J}$.
Denote by $\mathcal{T}(J)$ (resp., $\mathcal{F}(J)$) the category of $J$-torsion modules (resp., $J$-torsion free modules). It is clear from definitions that $\mathcal{T}(J)$ is closed under taking quotients: if $V$ lies in $\mathcal{T}(J)$ and $W$ is a submodule of $V$ in $\mathcal{O}\operatorname{-Mod}$, then $V/W$ lies in $\mathcal{T}(J)$ as well. Dually, $\mathcal{F}(J)$ is closed under taking submodules.
Given a rule $J$, we define its inclusion closure $\tilde{J}$ as follows: for every $x\in\operatorname{Ob}({\mathscr{C}})$,
$\tilde{J}(x)=\{S\subseteq{\mathscr{C}}(x,-)\mid\,\exists\,T\in J(x)\text{ suchthat }T\subseteq S\}.$
Clearly, $J$ is closed under inclusions if and only if $\tilde{J}=J$.
Lemma 3.4.
The inclusion closure $\tilde{J}$ satisfies the stability axiom, and $\mathcal{T}(J)=\mathcal{T}(\tilde{J})$.
Proof.
For the first statement, we take an arbitrary $x\in\operatorname{Ob}({\mathscr{C}})$, an arbitrary $S\in\tilde{J}(x)$, and an arbitrary morphism $f:x\to y$ in ${\mathscr{C}}$. By definition, there is a certain $T\in J(x)$ such that $T\subseteq S$. Then one has $f^{\ast}(S)\supseteq f^{\ast}(T)$. But $f^{\ast}(T)\in J(y)$ since $J$ satisfies the stability axiom. By the definition of $\tilde{J}$, one shall have $f^{\ast}(S)\in\tilde{J}(y)$, so $\tilde{J}$ also satisfies the stability axiom.
To establish the second statement, one notes that $J$-torsion modules are generated by $J$-torsion elements. Therefore, it suffices to show that $J$-torsion elements coincide with $\tilde{J}$-torsion elements for any $\mathcal{O}$-module $V$. Since every $S\in J(x)$ is clearly contained in $\tilde{J}(x)$ for $x\in\operatorname{Ob}({\mathscr{C}})$, $J$-torsion elements are clearly $\tilde{J}$-torsion elements. On the other hand, if an element $v$ in a certain $V_{x}$ is a $\tilde{J}$-torsion element, then one can find some $S\in\tilde{J}(x)$ such that $S\cdot v=0$. But by the definition of $\tilde{J}$, there exists some $T\in J(x)$ such that $T\subseteq S$, and in particular $T\cdot v=0$, so $v$ is also $J$-torsion. ∎
Remark 3.5.
By this lemma, when considering $J$-torsion modules, without loss of generality we can always assume that $J$ is closed under inclusions.
When $\mathcal{O}$ is a constant structure sheaf given by a ring which is still denoted by $\mathcal{O}$, for $S\in J(x)$, we have a short exact sequence of $\mathcal{O}$-modules
$0\to\underline{S}\to P(x)\to P(x)/\underline{S}\to 0$
where $P(x)$ and $\underline{S}$ are the $\mathcal{O}$-linearizations of ${\mathscr{C}}(x,-)$ and $S$, respectively. We need to generalize this construction to all structure sheaves as the quotient $\mathcal{O}$-module in the above sequence will play a key role for us to prove many results.
Denote by ${\mathscr{C}}_{x}$ the subcategory with the single object $x$ and the identity morphism $1_{x}$. The inclusion functor $\iota_{x}:{\mathscr{C}}_{x}\to{\mathscr{C}}$ induces a restriction functor $\textsl{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{x}\operatorname{-Mod}$ and its left adjoint $\textsl{ind}:\mathcal{O}_{x}\operatorname{-Mod}\to\mathcal{O}\operatorname{-Mod}$ is given by the left Kan extension. Explicitly, given an $\mathcal{O}_{x}$-module $W$ as well as an object $y$ in ${\mathscr{C}}$, one has
[…]
where $\mathcal{O}_{y}$ is regarded as a right $\mathcal{O}_{x}$-module for each $f$ via the ring homomorphism $\mathcal{O}_{f}:\mathcal{O}_{x}\to\mathcal{O}_{y}$. To deduce the second equility, we note that the colimit is taken over the category whose objects are morphisms $f:x\to y$ in ${\mathscr{C}}$ and whose morphisms are morphisms $h$ in ${\mathscr{C}}_{x}$ making the following diagram commutes
${x\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{h}$${y.}$${x\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$
It is easy to see that this is a discrete category since ${\mathscr{C}}_{x}$ only has the single morphism $1_{x}$, and consequently the colimit is actually a coproduct.
Now let $P(x)=\textsl{ind}(\mathcal{O}_{x})$, and $\underline{S}$ be the submodule such that
$\underline{S}_{y}=\bigoplus_{(f:\,x\to y)\in S}(\mathcal{O}_{y}\otimes_{\mathcal{O}_{x}}\mathcal{O}_{x})\cong\bigoplus_{(f:\,x\to y)\in S}\mathcal{O}_{y,f}\subseteq\bigoplus_{f\in{\mathscr{C}}(x,y)}\mathcal{O}_{y,f}=P(x)_{y}$
where $\mathcal{O}_{y,f}=\mathcal{O}_{y}$ for each $f$. It is easy to see that $P(x)/\underline{S}=0$ if and only if $S={\mathscr{C}}(x,-)$. In the case that $S\neq{\mathscr{C}}(x,-)$, the value $(P(x)/\underline{S})_{x}$ of $P(x)/\underline{S}$ on $x$ is a nonzero free $\mathcal{O}_{x}$-module since it contains a direct summand $\mathcal{O}_{x,1_{x}}$. Denote the identity in this direct summand by $\boldsymbol{1}_{x}$.
Lemma 3.6.
Given two rules $J$ and $K$, $\mathcal{T}(K)$ is a full subcategory of $\mathcal{T}(J)$ if and only if $K\subseteq\tilde{J}$.
Proof.
If $K\subseteq\tilde{J}$, then by the definition $\mathcal{T}(K)$ is a full subcategory of $\mathcal{T}(\tilde{J})$, which coincides with $\mathcal{T}(J)$ by Lemma 3.4.
Now we prove the only if direction. That is, for an arbitrary $x\in\operatorname{Ob}({\mathscr{C}})$ and $S\in K(x)$, we want to show that $S\in\tilde{J}(x)$. The conclusion holds trivially for $S={\mathscr{C}}(x,-)$, so we assume that $S$ is not ${\mathscr{C}}(x,-)$. Note that a morphism $f:x\to y$ in $S$ induces a homomorphism of $\mathcal{O}_{x}$-modules from $\mathcal{O}_{x,1_{x}}\subseteq P(x)_{x}$ to $\mathcal{O}_{y,f}\subseteq P(x)_{y}$. Passing to the quotient, we conclude that the identity $\boldsymbol{1}_{x}$ of $\mathcal{O}_{x,1_{x}}$ in $(P(x)/\underline{S})_{x}$ is a $K$-torsion element since $f\cdot\boldsymbol{1}_{x}=0$ for any $f\in S$. Thus $P(x)/\underline{S}$ is a $K$-torsion module since it is generated by $\boldsymbol{1}_{x}$ as an $\mathcal{O}$-module, so it is also a $J$-torsion module as $\mathcal{T}(K)$ is a full subcategory of $\mathcal{T}(J)$.
Although $\boldsymbol{1}_{x}$ might not be a $J$-torsion element, we can write $\boldsymbol{1}_{x}=e_{1}+\ldots+e_{r}$ where each $e_{i}$ is a nonzero $J$-torsion element. In particular, for $i=1$, we can find a certain $S_{1}\in J(x)$ such that $f\cdot e_{1}=0$ in $P(x)/\underline{S}$ for each $(f:x\to\bullet)\in S_{1}$. Since $f$ sends $e_{1}$ to $\boldsymbol{1}_{\bullet}\otimes e_{1}\in\mathcal{O}_{\bullet,f}\subseteq P(x)_{\bullet}$, the image of $\boldsymbol{1}_{\bullet}\otimes e_{1}$ in the quotient $(P(x)/\underline{S})_{\bullet}$ is 0 if and only if $f\in S$. Thus we have $f\in S$, and hence $S_{1}\subseteq S$. It follows that $S\in\tilde{J}(x)$ as desired. ∎
Remark 3.7.
We shall point out that the above proof works for the special case that $S=\emptyset$. In this situation, $\underline{S}$ is the zero module, so $P(x)=P(x)/\underline{S}$, and $S_{1}=\emptyset=S$.
Proposition 3.8.
Let $J$ and $K$ be two rules. Then $\tilde{J}=\tilde{K}$ if and only if $\mathcal{T}(J)=\mathcal{T}(K)$. In particular, if $J$ and $K$ are Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$, then $J=K$ if and only if $\mathcal{T}(J)=\mathcal{T}(K)$.
Proof.
The only if direction follows from Lemma 3.4. For the other direction, note that by Lemma 3.6 one shall have $J\subseteq\tilde{K}$ and $K\subseteq\tilde{J}$, so $\tilde{J}\subseteq\tilde{K}$ and $\tilde{K}\subseteq\tilde{J}$. ∎
We have described a procedure to construct a full subcategory $\mathcal{T}(J)$ of $\mathcal{O}\operatorname{-Mod}$ for a rule $J$. Now let us consider an inverse procedure.
Definition 3.9.
Given an $\mathcal{O}$-module $V$, an object $x\in\operatorname{Ob}({\mathscr{C}})$, and an element $v\in V_{x}$, we define the annihilator of $v$ to be
$\mathcal{A}(v)=\{f:x\to\bullet\mid f\cdot v=0\},$
[…]
$\mathcal{A}(\mathfrak{A})(x)=\{\mathcal{A}(v)\mid v\in V_{x},\,V\in\operatorname{Ob}(\mathfrak{A})\}.$
[…]
Lemma 3.10.
[…]
Proof.
Take $S\in\mathcal{A}(\mathfrak{A})(x)$ and an arbitrary morphism $f:x\to y$. If $S\neq\emptyset$, then there is an $\mathcal{O}$-module $V$ in $\mathfrak{A}$ as well as an element $v\in V_{x}$ such that $S=\mathcal{A}(v)$. For $f\cdot v\in V_{y}$, we have
$f^{\ast}(S)=\{g:y\to\bullet\mid gf\in S\}=\{g:y\to\bullet\mid g\cdot(f\cdot v)=0\}=\mathcal{A}(f\cdot v).$
Thus $f^{\ast}(S)\in\mathcal{A}(\mathfrak{A})(y)$. For $S=\emptyset$, we can find an element $v\in V_{x}$ such that $f\cdot v\neq 0$ for all $f\in{\mathscr{C}}(x,-)$. In particular, for a fixed $f:x\to y$, we have
$\emptyset=f^{\ast}(S)=\{g:y\to\bullet\mid g\cdot(f\cdot v)=(gf)\cdot v=0\}=\mathcal{A}(f\cdot v)\in\mathcal{A}(\mathfrak{A})(y),$
so the stability axiom also holds for this case. ∎
The following result can be viewed as an analogue of the well known correspondence between algebraic sets and radical ideals given by Hilbert’s Nullstellensatz. In particular, $\mathcal{T}$ is the analogue of the operation taking algebraic sets, and $\mathcal{A}$ is the analogue of the operation taking ideals associated to algebraic sets. With respect to these analogues, the full subcategory of $J$-torsion $\mathcal{O}$-modules, where $J$ is closed under finite intersections, plays the role of an algebraic set, and the rule $J$ closed under finite intersections and inclusions plays the role of a radical ideal.
Proposition 3.11.
Given a rule $J$, one has:
  1. (1)
    $\mathcal{T}(\mathcal{A}(\mathcal{T}(J)))=\mathcal{T}(J)$ if $J$ is closed under finite intersections;
  2. (2)
    $J=\mathcal{A}(\mathcal{T}(J))$ if and only if $J$ is closed under finite intersections and inclusions.
Proof.
For brevity denote $\mathcal{A}(\mathcal{T}(J))$ by $K$.
(1): For $x\in\operatorname{Ob}({\mathscr{C}})$ and $S\in J(x)$, consider the $\mathcal{O}$-module $P(x)/\underline{S}$ defined before Lemma 3.6. It is a $J$-torsion module, and hence is contained in $\mathcal{T}(J)$. By the definition of $K$, we conclude that $S=\mathcal{A}(\boldsymbol{1}_{x})\in K(x)$, so $J\subseteq K$. Consequently, $\mathcal{T}(J)$ is a full subcategory of $\mathcal{T}(K)$. Now we prove $\mathcal{T}(K)\subseteq\mathcal{T}(J)$ by showing that every $K$-torsion $\mathcal{O}$-module $V$ is also $J$-torsion. Since torsion modules are generated by torsion elements, it suffices to show the following claim: for $x\in\operatorname{Ob}({\mathscr{C}})$ and $v\in V_{x}$, if $v$ is $K$-torsion, then it is $J$-torsion.
Assuming that $v$ is $K$-torsion, we can find a certain $S\in K(x)$ such that $S\cdot v=0$. By definition of $K$, we can find a $J$-torsion $\mathcal{O}$-module $W$ and an element $w\in W_{x}$ such that $S=\mathcal{A}(w)$. But $J$ is closed under finite intersections, so $w$ is a $J$-torsion element by Lemma 3.2, and hence there is a certain $S^{\prime}\in J(x)$ such that $S^{\prime}\cdot w=0$. Clearly, we shall have $S^{\prime}\subseteq S=\mathcal{A}(w)$, so $S^{\prime}\cdot v=0$. Consequently, $v$ is a $J$-torsion element as well, and the claim is established.
(2): We have shown that $J\subseteq K$ in the proof of (1), so it suffices to prove $K\subseteq J$ to establish the if direction. The given assumption tells us that $J=\tilde{J}$. Thus by Lemma 3.6, it suffices to show that $\mathcal{T}(K)$ is a full subcategory of $\mathcal{T}(J)$. But this is true by (1).
Now we turn to the only if direction. One first checks that $J=K$ is closed under finite intersections. For $x\in\operatorname{Ob}({\mathscr{C}})$ and $S,S^{\prime}\in K(x)$, we can find objects $V$ and $V^{\prime}$ in $\mathcal{T}(J)$ and elements $v\in V_{x}$ and $v^{\prime}\in V^{\prime}_{x}$ such that $S=\mathcal{A}(v)$ and $S^{\prime}=\mathcal{A}(v^{\prime})$. Note that $\mathcal{T}(J)$, though might not be closed under extensions, is closed under finite direct sums. Consequently, $V\oplus V^{\prime}$ is contained in $\mathcal{T}(J)$. In particular, for $(v,v^{\prime})\in V_{x}\oplus V_{x}^{\prime}$, one has $S\cap S^{\prime}=\mathcal{A}((v,v^{\prime}))\in K(x)$, so $K$ is closed under finite intersections.
Finally we check that $J$ is closed under inclusions. Suppose that $S\subseteq T\subseteq{\mathscr{C}}(x,-)$ and $S\in J(x)$. Consider $\mathcal{O}$-modules $P(x)/\underline{S}$ and $P(x)/\underline{T}$. Note that $P(x)/\underline{S}$ is contained in $\mathcal{T}(J)$, $P(x)/\underline{T}$ is a quotient of $P(x)/\underline{S}$, and $\mathcal{T}(J)$ is closed under taking quotients. Consequently, $P(x)/\underline{T}$ is also contained in $\mathcal{T}(J)$. Thus $T=\mathcal{A}(\boldsymbol{1}_{x})\in K(x)=J(x)$, so $J$ is closed under inclusions. ∎
The above proof shows that $\mathcal{A}(\mathcal{T}(J))$ is always closed under finite intersections.
Remark 3.12.
We illustrate the above result from a categorical viewpoint. Let $\mathcal{J}$ be the set of rules $J$ satisfying the stability axiom. It is a poset under the inclusion. We can construct another category whose objects are full subcategories of $\mathcal{O}\operatorname{-Mod}$ and morphisms are inclusion functors. Then $\mathcal{A}$ and $\mathcal{T}$ are covariant functors between these two categories. Restricting to the subposet consisting of $J$ closed under finite intersections, we have $\mathcal{T}\mathcal{A}\mathcal{T}=\mathcal{T}$; and restricting further to the subposet consisting of $J$ closed under finite intersections and inclusions, $\mathcal{A}\mathcal{T}$ is the identity functor.
It is not hard to see that $\mathcal{T}$ preserves finite intersections. Explicitly, given two rules $J$ and $K$ closed under inclusions and finite intersections, one has $\mathcal{T}(J)\cap\mathcal{T}(K)=\mathcal{T}(J\cap K)$, where $J\cap K$ is the rule assigning $J(x)\cap K(x)$ to $x\in\operatorname{Ob}({\mathscr{C}})$. Indeed, since $\mathcal{T}(J\cap K)\subseteq\mathcal{T}(J)$ and $\mathcal{T}(J\cap K)\subseteq\mathcal{T}(K)$, one has $\mathcal{T}(J)\cap\mathcal{T}(K)\supseteq\mathcal{T}(J\cap K)$. On the other hand, take $V$ in $\mathcal{T}(J)\cap\mathcal{T}(K)$. Then for $x\in\operatorname{Ob}({\mathscr{C}})$ and $v\in V_{x}$, one can find $S\in J(x)$ and $T\in K(x)$ such that $S\cdot v=0=T\cdot v$. Then $S\cup T\in J(x)\cap K(x)$ and $(S\cup T)\cdot v=0$, so $v$ is also $J\cap K$-torsion, and hence $V\in\mathcal{T}(J\cap K)$.
3.2. Torsion pairs induced by Grothendieck topologies
In this subsection we show that Grothendieck topologies induce hereditary torsion pairs in $\mathcal{O}\operatorname{-Mod}$, and are completely determined by the induced torsion pairs.
Proposition 3.13.
Given a rule $J$, one has:
  1. (1)
    $(\mathcal{T}(J),\,\mathcal{F}(J))$ is a hereditary torsion pair in $\mathcal{O}\operatorname{-Mod}$ if $J$ is a Grothendieck topology.
  2. (2)
    Conversely, if $J$ is closed under inclusions and $(\mathcal{T}(J),\,\mathcal{F}(J))$ is a hereditary torsion pair in $\mathcal{O}\operatorname{-Mod}$, then $J$ is a Grothendieck topology.
Proof.
(1): Firstly, $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,F)=0$ for $T$ in $\mathcal{T}(J)$ and $F$ in $\mathcal{F}(J)$. Indeed, take any morphism $\phi:T\to F$. By (2) of Lemma 3.2, $\phi$ restricts to a morphism
$\mathcal{T}_{J}(T)=T\longrightarrow\mathcal{T}_{J}(F)=0,$
which is the zero morphism. Therefore, $\phi=0$. Secondly, it is clear that $\mathcal{T}(J)$ is closed under taking quotient modules. It is also closed under taking submodules. Indeed, for any $V$ in $\mathcal{T}(J)$, $x\in\operatorname{Ob}({\mathscr{C}})$ and $v\in V_{x}$, $v$ is a $J$-torsion element by Lemma 3.2. Consequently, for an arbitrary submodule $W\subseteq V$ and every $x\in\operatorname{Ob}({\mathscr{C}})$, all elements in $W_{x}$ are also $J$-torsion.
Clearly, only the zero module lies in the intersection of $\mathcal{T}(J)$ and $\mathcal{F}(J)$. To deduce the conclusion, it suffices to show that $V/\mathcal{T}_{J}(V)$ is $J$-torsion free for every $\mathcal{O}$-module $V$; that is, for $x\in\operatorname{Ob}({\mathscr{C}})$ and $\bar{v}\in(V/\mathcal{T}_{J}(V))_{x}$, if $\bar{v}$ is $J$-torsion, then $\bar{v}=0$.
If $\emptyset\in J(x)$, then every $v\in V_{x}$ is $J$-torsion, so the conclusion holds trivially as $(V/\mathcal{T}_{J}(V))_{x}=0$. Otherwise, take a preimage $v\in V_{x}$ of $\bar{v}$. Since $\bar{v}$ is $J$-torsion, there exists a nonempty $S\in J(x)$ such that $f\cdot v\in(\mathcal{T}_{J}(V))_{\bullet}$ for each $f:x\to\bullet$ in $S$. But $f\cdot v$ is $J$-torsion by Lemma 3.2, so we can find a certain $S_{f}\in J(\bullet)$ such that $S_{f}\cdot(f\cdot v)=0$. Define
$\hat{S}=\bigcup_{f\in S}S_{f}f=\bigcup_{f\in S}\{gf\mid g\in S_{f}\},$
which is clearly a subfunctor of ${\mathscr{C}}(x,-)$ and $\hat{S}\cdot v=0$. Since for every $f\in S$,
$f^{\ast}(\hat{S})\supseteq f^{\ast}(S_{f}f)\supseteq S_{f}\in J(\bullet),$
one has $f^{\ast}(\hat{S})\in J(\bullet)$ because $J$ is closed under inclusions, and hence $\hat{S}\in J(x)$ by the transitivity axiom. Consequently, $v$ is $J$-torsion and is contained in $(\mathcal{T}_{J}(V))_{x}$, so $\bar{v}=0$ as desired.
(2) The maximal axiom holds trivially as $J$ is closed under inclusions. Since we have assumed that $J$ always satisfies the stability axiom, it remains to check the transitivity axiom. Given $T\in J(x)$ and $S\subseteq{\mathscr{C}}(x,-)$ such that $g^{\ast}(S)\in J(y)$ for every $g:x\to y$ in $T$, we need to show that $S\in J(x)$. Of course, we can assume $S\neq{\mathscr{C}}(x,-)$.
The torsion pair gives a short exact sequence
$0\to P^{\prime}\to P(x)/\underline{S}\to P^{\prime\prime}\to 0$
where $P(x)/\underline{S}$ is constructed before Lemma 3.6, $P^{\prime}$ is $J$-torsion and $P^{\prime\prime}$ is $J$-torsion free. Consider $\boldsymbol{1}_{x}\in\mathcal{O}_{x,1_{x}}\subseteq(P(x)/\underline{S})_{x}$ which is specified before Lemma 3.6. For every $g:x\to y$ in $T$, since $g^{\ast}(S)\in J(y)$ and $g^{\ast}(S)\cdot(g\cdot\boldsymbol{1}_{x})=0$, it follows that $g\cdot\boldsymbol{1}_{x}$ is $J$-torsion, and hence is contained in $P^{\prime}_{y}$. Passing to the quotient module $P^{\prime\prime}$, we deduce that $T\cdot\bar{\boldsymbol{1}}_{x}=0$, where $\bar{\boldsymbol{1}}_{x}$ is the image of $\boldsymbol{1}_{x}$ in $P_{x}^{\prime\prime}$. Thus $\bar{\boldsymbol{1}}_{x}$ is $J$-torsion. But $P^{\prime\prime}$ is torsion free, this forces $\bar{\boldsymbol{1}}_{x}=0$, or equivalently $\boldsymbol{1}_{x}$ is contained in $P^{\prime}_{x}$. But $P(x)/\underline{S}$ as an $\mathcal{O}$-module is generated by $\boldsymbol{1}_{x}$, this forces $P(x)/\underline{S}=P^{\prime}$; that is, $P(x)/\underline{S}$ is a $J$-torsion module. Now we can apply the same argument as in the last paragraph of the proof of Lemma 3.6 to deduce $S\in J(x)$. ∎
Remark 3.14.
If $(\mathcal{T}(J),\,\mathcal{F}(J))$ is a hereditary torsion pair, then $\mathcal{T}(J)$ is a Serre subcategory of $\mathcal{O}\operatorname{-Mod}$, so we can form a Serre quotient $\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J)$. In the next section we will show that it is equivalent to ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ when $\mathcal{O}$ is a structure sheaf.
Remark 3.15.
The condition that $J$ is closed under inclusions in part (2) of the previous proposition cannot be dropped. For instance, take $J$ to be the maximal topology on ${\mathscr{C}^{\mathrm{op}}}$, and let $K$ be the rule such that $K(x)=\{\emptyset,\,{\mathscr{C}}(x,-)\}$ for each $x\in\operatorname{Ob}({\mathscr{C}})$. It is easy to see that $K$ satisfies the stability axiom and the maximal axiom. Furthermore, $J$ and $K$ induce the same hereditary torsion pair on $\mathcal{O}\operatorname{-Mod}$ with $\mathcal{T}(J)=\mathcal{T}(K)=\mathcal{O}\operatorname{-Mod}$ and $\mathcal{F}(J)=\mathcal{F}(K)=\{0\}$. But in general $K$ is not a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$.
We end this section by an example illustrating our constructions.
Example 3.16.
Let ${\mathscr{C}}$ be the free category of the following quiver
${x\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{g}$${y.}$
By a direct computation, we obtain all possible Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$, listed below:
Topologies $J(x)$ $J(y)$
Trivial topology $\{\{1_{x},\,f,\,g\}\}$$\{\{1_{y}\}\}$
Dense topology $\{\{1_{x},\,f,\,g\},\,\{f,\,g\}\}$$\{\{1_{y}\}\}$
Maximal topology $\{\{1_{x},\,f,\,g\},\{f,\,g\},\,\{f\},\,\{g\},\,\emptyset\}$$\{\{1_{y}\},\,\emptyset\}$
Topology IV $\{\{1_{x},\,f,\,g\}\}$$\{\{1_{y}\},\,\emptyset\}$.
The corresponded torsion pairs induced by these Grothendieck topologies are:
Topologies $\mathcal{T}(J)$ $\mathcal{F}(J)$
Trivial topology 00 $\mathcal{O}\operatorname{-Mod}$
Dense topology $\{V\in\mathcal{O}\operatorname{-Mod}\mid V_{y}=0\}$$\{V\in\mathcal{O}\operatorname{-Mod}\mid\ker V_{f}\cap\ker V_{g}=0\}$
Maximal topology $\mathcal{O}\operatorname{-Mod}$ 00
Topology IV $\{V\in\mathcal{O}\operatorname{-Mod}\mid V_{x}=0\}$$\{V\in\mathcal{O}\operatorname{-Mod}\mid V_{y}=0\}$.
4. A torsion theoretic characterization of sheaves
In this section we interpret notions in sheaf theory in terms of the torsion theory developed in the previous section. We mention that all results have been proved in [7] for the special case that $J$ is the atomic Grothendieck topology and $\mathcal{O}$ is a constant structure sheaf. Furthermore, the strategy of proofs in that paper extend to the general framework in this paper with suitable modifications.
Throughout this subsection let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ and let $\mathcal{O}$ be a structure sheaf on the site $({\mathscr{C}^{\mathrm{op}}},J)$. Recall that an $\mathcal{O}$-module $V$ is right perpendicular to all $J$-torsion $\mathcal{O}$-modules in the sense of Geigle and Lenzing [12] if $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,V)=0=\operatorname{Ext}_{\mathcal{O}\operatorname{-Mod}}^{1}(T,V)$ for each $J$-torsion module $T$.
Lemma 4.1.
Let $V$ be an $\mathcal{O}$-module. The following statements are equivalent:
  1. (1)
    $V$ is a $J$-torsion free $\mathcal{O}$-module;
  2. (2)
    $V$ is a separated presheaf of modules;
  3. (3)
    $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,V)=0$ for $T\in\mathcal{T}(J)$.
Proof.
It is clear that $V$ is $J$-torsion free if and only if (3) holds, so we only need to show the equivalence of (1) and (2). If $V$ is not $J$-torsion free, then there exist $x\in\operatorname{Ob}({\mathscr{C}})$, $0\neq v\in V_{x}$, and $S\in J(x)$ such that $S\cdot v=0$. Now by setting $v_{f}=0\in V_{\bullet}$ for each $(f:x\to\bullet)\in S$, we get a matching family which has at least two amalgamations, namely $0\in V_{x}$ and $v\in V_{x}$. Thus $V$ is not separated. Conversely, if $V$ is not separated, we can find $x\in\operatorname{Ob}({\mathscr{C}})$, $S\in J(x)$, and a matching family $\{v_{f}\in V_{\bullet}\mid(f:x\to\bullet)\in S\}$ with two distinct amalgamations $v,w\in V_{x}$. But in this case $v-w$ is sent to 0 by any $f\in S$, so $V$ is not $J$-torsion free. This finishes the proof. ∎
Now we are ready to describe homological characterizations of sheaves of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$.
Theorem 4.2.
Let $V$ be an $\mathcal{O}$-module. The following statements are equivalent:
  1. (1)
    $V$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$;
  2. (2)
    $V$ is a $J$-saturated $\mathcal{O}$-module;
  3. (3)
    $V$ is right perpendicular to all $J$-torsion modules.
Proof.
By Lemma 4.1, to prove the theorem, we can assume that $V$ is a $J$-torsion free $\mathcal{O}$-module.
$(2)\Leftrightarrow(3)$. In [7, Lemma 3.6] we have proved the conclusion for the special case that $J$ is the atomic Grothendieck topology. We use the same strategy to establish the equivalence for all Grothendieck topologies.
Suppose that $V$ is $J$-torsion free. It has an injective hull $E$ in $\mathcal{O}\operatorname{-Mod}$ since $\mathcal{O}\operatorname{-Mod}$ is a Grothendieck category. We claim that $E$ is torsion free. Otherwise, we have $\mathcal{T}_{J}(V)\neq 0$. But $V\cap\mathcal{T}_{J}(V)=0$ since $V$ is $J$-torsion free. This is impossible because the inclusion $V\to E$ is essential by [23, Definition 47.2.1]. Thus $E$ is $J$-torsion free and injective. Applying $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(T,-)$ to the short exact sequence $0\to V\to E\to E/V\to 0$ of $\mathcal{O}$-modules we deduce that $V$ is right perpendicular to every $J$-torsion module $T$ if and only if $E/V$ is $J$-torsion free. But this is equivalent to the condition that $V$ is $J$-saturated via applying the functor $\mathcal{T}_{J}$ to the short exact sequence.
$(1)\Rightarrow(2)$. Let $V$ be a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$. It is $J$-torsion free. Since the category of $J$-torsion modules is a localizing subcategory by [12, Proposition 2.5], applying [12, Proposition 2.2] we obtain a short exact sequence
$0\to V\to\tilde{V}\to W\to 0$
such that $W$ is $J$-torsion and $\tilde{V}$ is right perpendicular to all $J$-torsion modules, and hence $J$-saturated by the equivalence of (2) and (3). To show that $V$ is $J$-saturated, it suffices to check that $W$ is $J$-torsion free, which forces $W=0$ and hence $V\cong\tilde{V}$.
Let $w\in W_{x}$ be a $J$-torsion element for some $x\in\operatorname{Ob}({\mathscr{C}})$. We can find a certain $S\in J(x)$ such that $S\cdot w=0$. Let $\tilde{w}\in\tilde{V}_{x}$ be a preimage of $w$, so we have $f\cdot\tilde{w}\in V_{y}$ for every $f:x\to y$ in $S$. It is clear that the rule
$(f:x\to y)\in S\mapsto f\cdot\tilde{w}$
is a matching family for $S$ of elements in $V$. Since $V$ is a sheaf of modules, we can find an amalgamation $\tilde{u}\in V_{x}$ such that $f\cdot\tilde{u}=f\cdot\tilde{w}$ for every $f\in S$. We can view $\tilde{u}$ as an element in $\tilde{V}_{x}$ by identifying $V$ with a submodule of $\tilde{V}$. Thus in the $\mathcal{O}$-module $\tilde{V}$ one has $f\cdot(\tilde{u}-\tilde{w})=0$ for $f\in S$. But $\tilde{V}$ is $J$-torsion free, so $\tilde{w}=\tilde{u}\in V_{x}$, and hence $w=0$. Consequently, $W$ is $J$-torsion free.
$(2)\Rightarrow(1)$: Suppose that $V$ is $J$-saturated. It is a separated presheaf of modules by Lemma 4.1. Take an arbitrary $x\in\operatorname{Ob}({\mathscr{C}})$, a certain $S\in J(x)$, and a matching family for $S$ of elements in $V$
$\{v_{f}\in V_{y}\mid(f:x\to y)\in S\}.\quad\quad({\dagger})$
We need to show the existence of amalgamations. By [18, III.5, Lemma 2], the canonical map $\eta_{V}:V\to V^{+}$ is injective where $V^{+}=V^{\sharp}$ is the sheafification of $V$. Thus we obtain the following short exact sequence in $\mathcal{O}\operatorname{-Mod}$:
$0\to V\to V^{+}\to W\to 0$
Furthermore, by [18, III.5, Lemma 5], $V^{+}$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, and hence a $J$-saturated $\mathcal{O}$-module by the conclusion of the only if direction. Applying $\mathcal{T}_{J}$ to this short exact sequence, we deduce that $W$ is $J$-torsion free since $V$ is $J$-saturated.
[…]
We describe a few consequences of the above theorem. The first one can be viewed as a ringed version of Remark 2.6.
Corollary 4.3.
Given an $\mathcal{O}$-module $V$, the following statements are equivalent:
  1. (1)
    $V$is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$;
  2. (2)
    for any $x\in\operatorname{Ob}({\mathscr{C}})$ and any $S\in J(x)$, the inclusion $\underline{S}\to P(x)$ induces an isomorphism
    $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(P(x),V)\cong\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(\underline{S},V),$
    where $P(x)$ and $\underline{S}$ are as defined before Lemma 3.6;
  3. (3)
    for any $x\in\operatorname{Ob}({\mathscr{C}})$ and any $S\in J(x)$, the inclusion $\mathbb{Z}S\to\mathbb{Z}{\mathscr{C}}(x,-)$ induces an isomorphism
    $\operatorname{Hom}_{\underline{\mathbb{Z}}\operatorname{-Mod}}(\mathbb{Z}{\mathscr{C}}(x,-),V)\cong\operatorname{Hom}_{\underline{\mathbb{Z}}\operatorname{-Mod}}(\mathbb{Z}S,V),$
    where $\underline{\mathbb{Z}}$ is the constant structure sheaf induced by the ring of integers and $\mathbb{Z}S$ is the free $\mathbb{Z}$-module spanned by elements in $S$.
Proof.
(1$\Rightarrow$(2). Applying $\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(-,V)$ to the short exact sequence
$0\to\underline{S}\to P(x)\to P(x)/\underline{S}\to 0$
we obtain another exact sequence
$0\to\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(P(x)/\underline{S},V)\to\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(P(x),V)$
$\to\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(\underline{S},V)\to\operatorname{Ext}^{1}_{\mathcal{O}\operatorname{-Mod}}(P(x)/\underline{S},V)\to 0$
by noting that $P(x)$ is projective in $\mathcal{O}\operatorname{-Mod}$ (this is clear when $\mathcal{O}$ is a constant structure sheaf, and it actually holds for arbitrary structure sheaves). Since $V$ is right perpendicular to all $J$-torsion $\mathcal{O}$-modules by the previous theorem and $P(x)/\underline{S}$ is a $J$-torsion module, the first term and the last term vanish. Thus we have the desired isomorphism.
(2$\Rightarrow$(1). By the previous theorem, it suffices to show that
$\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(U,V)=0=\operatorname{Ext}^{1}_{\mathcal{O}\operatorname{-Mod}}(U,V)$
for any $J$-torsion module $U$. The long exact sequence in the previous paragraph and the given isomorphism tell us that these identities hold for $U=P(x)/\underline{S}$ with $x\in\operatorname{Ob}({\mathscr{C}})$ and $S\in J(x)$. But
$\mathcal{P}=\{P(x)/\underline{S}\mid x\in\operatorname{Ob}({\mathscr{C}}),\,S\in J(x)\}$
is a set of generators of $\mathcal{T}(J)$, so we may construct a short exact sequence
$0\to U^{\prime}\to Q\to U\to 0$
where $Q$ is a coproduct of $\mathcal{O}$-modules in $\mathcal{P}$. Consider the long exact sequence
$0\to\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(U,V)\to\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(Q,V)\to\operatorname{Hom}_{\mathcal{O}\operatorname{-Mod}}(U^{\prime},V)$
$\to\operatorname{Ext}^{1}_{\mathcal{O}\operatorname{-Mod}}(U,V)\to\operatorname{Ext}^{1}_{\mathcal{O}\operatorname{-Mod}}(Q,V).$
The second and the fifth terms vanish by the structure of $Q$. It follows that the first term vanishes as well. But since $U$ can be any object in $\mathcal{T}(J)$ and $U^{\prime}$ is contained in $\mathcal{T}(J)$, the third term also vanishes. Consequently, $\operatorname{Ext}^{1}_{\mathcal{O}\operatorname{-Mod}}(U,V)=0$, and the proof is complete.
(1$\Leftrightarrow$(3). Note that $V$ is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ if and only if it is a sheaf of abelian groups; that is, viewed as a $\underline{\mathbb{Z}}$-module it is a sheaf of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\underline{\mathbb{Z}})$. The equivalence between (1) and (3) then follows from this observation as well as the equivalence between (1) and (2). ∎
The second one gives a characterization of injective sheaves.
Corollary 4.4.
Injective objects in ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ coincide with injective $J$-torsion free $\mathcal{O}$-modules.
Proof.
This result has been proved in [7, Corollary 3.9] for the atomic Grothendieck topology. Its proof still works in full generality. ∎
Let $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$ be the full subcategory of $\mathcal{O}\operatorname{-Mod}$ consisting of $J$-saturated $\mathcal{O}$-modules. The third consequence of Theorem 4.2 provides the following equivalence:
Corollary 4.5.
One has the following identification and equivalence of categories:
[…]
Proof.
The identification follows from Theorem 4.2  [^3] We shall remind the reader that the abelian structure on $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$ is not inherited from the abelian structure on $\mathcal{O}\operatorname{-Mod}$. Indeed, Given a monomorphism $\alpha:V\to W$ between two $J$-saturated modules, its cokernel in $\mathcal{O}\operatorname{-Mod}$ might not coincide with the cokernel in $\mathcal{O}\operatorname{-Mod}^{{J\text{-}\mathrm{sat}}}$., while the equivalence follows from [12, Proposition 2.2(d)]. ∎
Theorem 4.2 also provides a new description of the sheafification functor $\sharp$, which is more elementary compared to the ones using local isomorphisms or dense monomorphisms. This description has been observed for the atomic Grothendieck topology in [7, Section 3], and here we extend it to all Grothendieck topologies. Explicitly, given an $\mathcal{O}$-module $V$, the first step is to take the $J$-torsion free part $V_{F}=V/\mathcal{T}_{J}(V)$ and embeds it into an injective hull $I$ (which is $J$-saturated) to get a short exact sequence $0\to V_{F}\to I\to C\to 0$ of $\mathcal{O}$-modules. In the second step, we construct the pull back of $I\to C$ and $C_{T}=\mathcal{T}_{J}(C)\to C$ to obtain the following commutative diagram such that all rows and columns are exact:
[…]
Then $V^{\sharp}\cong\widetilde{V}$. For more details, please refer to [7, Section 3].
Now we consider sheaf cohomology groups. Given an object $x\in\operatorname{Ob}({\mathscr{C}})$, we can define a functor
$\Gamma_{x}^{p}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{x}\operatorname{-Mod},\quad V\mapsto V_{x},$
which is exact. Restricting it to the full subcategory ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$, we get a functor $\Gamma_{x}$ which is only left exact since the abelian structure of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ is not inherited from that of $\mathcal{O}\operatorname{-Mod}$. Thus one can define $\mathrm{R}^{i}\Gamma_{x}(V)$, the $i$-th right derived functor of $\Gamma_{x}$, called the $i$-th sheaf cohomology group of $V$ at $x$. For details, please refer to [23, I.21.2].
Sheaf cohomology groups can be computed via injective resolutions. The following result relates them to right derived functors of $\mathcal{T}_{J}$.
Corollary 4.6.
For $i\geqslant 1$, one has the following isomorphism of functors
$\mathrm{R}^{i}\Gamma_{x}\cong\Gamma_{x}^{p}\circ(\mathrm{R}^{i+1}\mathcal{T}_{J})$
Proof.
It has been proved in [7, Theorem 3.16] for the special case that $J$ is the atomic Grothendieck topology and $\mathcal{O}$ is a constant structure sheaf. The proof actually works in the general situation. ∎
It is well known that morphisms between ringed sites induce geometric morphisms between categories of sheaves of modules. In the rest of this section we give an illustration of this fact via our torsion theoretic interpretation. Let $({\mathscr{C}^{\mathrm{op}}},\,J,\,\mathcal{O})$ and $({\mathscr{D}}^{{\mathrm{op}}},\,K,\,\mathcal{O}^{\prime})$ be two ringed sites. To simplify the technical issue we suppose that both $\mathcal{O}$ and $\mathcal{O}^{\prime}$ are the constant structure sheaf $\underline{\mathbb{Z}}$ induced by $\mathbb{Z}$. This simplification does not cause essential influence to the general case since a sheaf of modules on $({\mathscr{C}^{\mathrm{op}}},\,J,\,\mathcal{O})$ is an $\mathcal{O}$-module as well as a sheaf of abelian groups.
Recall that a functor $\phi:{\mathscr{C}^{\mathrm{op}}}\to{\mathscr{D}}^{{\mathrm{op}}}$ is cover-preserving if for every $x\in\operatorname{Ob}({\mathscr{C}})$ and every $S\in J(x)$, the sieve generated by $\phi(S)=\{\phi(f)\mid f\in S\}$ is contained in $K(\phi(x))$.
Proposition 4.7.
Suppose that ${\mathscr{C}^{\mathrm{op}}}$ and ${\mathscr{D}}^{{\mathrm{op}}}$ have finite limits and $\phi:{\mathscr{C}^{\mathrm{op}}}\to{\mathscr{D}}^{{\mathrm{op}}}$ is cover-preserving and preserves finite limits. Then $\phi$ induces a geometric morphism
$\phi^{\ast}:{\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,K,\,\underline{\mathbb{Z}})\to{\mathrm{Sh}}({\mathscr{C}^{\mathrm{op}}},\,J,\,\underline{\mathbb{Z}}).$
Proof.
This is essentially a combination of [13, C2.3, Lemma 2.3.3] and [13, C2.3, Corollary 2.3.4], so we only give a sketch of the proof. The functor $\phi$ induces a restriction functor
$\tilde{\phi}^{\ast}:{\mathrm{PSh}}({\mathscr{D}}^{{\mathrm{op}}},\,\underline{\mathbb{Z}})\to{\mathrm{PSh}}({\mathscr{C}^{\mathrm{op}}},\,\underline{\mathbb{Z}})$
between the two presheaf categories as well as a left adjoint
$\phi_{!}:{\mathrm{PSh}}({\mathscr{C}^{\mathrm{op}}},\,\underline{\mathbb{Z}})\to{\mathrm{PSh}}({\mathscr{D}}^{{\mathrm{op}}},\,\underline{\mathbb{Z}})$
given by the left Kan extension along $\phi$. Since $\phi$ preserves finite limits, $\phi_{!}$ is exact. Thus it suffices to show that $\tilde{\phi}^{\ast}$ restricts to a functor
$\phi^{\ast}:{\mathrm{Sh}}({\mathscr{D}}^{{\mathrm{op}}},\,K,\,\underline{\mathbb{Z}})\to{\mathrm{Sh}}({\mathscr{C}^{\mathrm{op}}},\,J,\,\underline{\mathbb{Z}}).$
In this case, it has an exact left adjoint given by the composite of
[…]
To complete the proof, we need to show that $\tilde{\phi^{\ast}}$ sends $K$-saturated modules to $J$-saturated modules. Since $(\phi_{!},\,\tilde{\phi}^{\ast})$ is an adjoint pair and $\phi_{!}$ is exact, it suffices to show that $\phi_{!}$ sends $J$-torsion modules to $K$-torsion modules, which happens if and only if $\tilde{\phi}^{\ast}$ sends $K$-torsion free modules to $J$-torsion free modules. This is clear. Indeed, suppose that $V$ is $K$-torsion free. If there exist $x\in\operatorname{Ob}({\mathscr{C}})$ and $v\in(\tilde{\phi}^{\ast}V)_{x}$ such that $v$ is $J$-torsion, then one can find a covering sieve $S\in J(x)$ such that $S\cdot v=0$. But this means that $\phi(f)$ sends $v\in V_{\phi(x)}$ to 0 for each $f\in S$. Consequently, $v$ is $K$-torsion since the covering sieve in $K(\phi(x))$ generated by $\phi(S)$ sends it to 0. This forces $v=0$, so $\tilde{\phi}^{\ast}V$ is $J$-torsion free. ∎
Remark 4.8.
In the above proposition the conditions that ${\mathscr{C}^{\mathrm{op}}}$ and ${\mathscr{D}}^{{\mathrm{op}}}$ have finite limits and $\phi$ preserves finite limits are used to show the exactness of the left adjoint $\phi_{!}$. Since $\phi_{!}$ is precisely the tensor functor $-\otimes_{\mathbb{Z}{\mathscr{D}}^{{\mathrm{op}}}}\mathbb{Z}{\mathscr{C}}^{{\mathrm{op}}}$, for categories without finite limits, one can use the following condition: the category algebra $\mathbb{Z}{\mathscr{C}}^{{\mathrm{op}}}$ is a right flat $\mathbb{Z}{\mathscr{D}}^{{\mathrm{op}}}$-module.
5. Grothendieck topologies on artinian EI categories
In the previous section we have characterized sheaves of modules on $(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ as $J$-saturated $\mathcal{O}$-modules and obtained the following equivalence
${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})\simeq\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J).$
The Serre quotient on the right hand side is still complicated, so we want to find a more explicit description for some special Grothendieck topologies $J$. When $J$ is a subcategory topology (or called rigid topology), this Serre quotient is equivalent to $\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$, where $\mathcal{O}_{\mathscr{D}}$ is the restriction of $\mathcal{O}$ to a full subcategory $\mathscr{D}$ of ${\mathscr{C}}$, [^4] Since $\mathcal{O}$ is a covariant functor from $\mathcal{C}$ to the category of associative rings, its shall be restricted to ${\mathscr{D}}$ rather than ${\mathscr{D}}^{{\mathrm{op}}}$. and hence becomes much more transparent. Furthermore, it has been pointed out in [2, IV.9] that all Grothendieck topologies on a finite category are rigid, and this result also holds for artinian posets (see [17, Theorem 2.12]).
The main goal of this section is to prove the above conclusion for EI categories whose underlying posets are artinian. In particular, we generalize the corresponding results for finite acyclic quivers by Murfet [19] and for posets by Hemelaer [14] and Lindenhovius [17].
5.1. Definitions and the main result
Given a skeletal small category ${\mathscr{C}}$, we impose the following relation $\leqslant$ on $\operatorname{Ob}({\mathscr{C}})$: $x\leqslant y$ if ${\mathscr{C}}(x,y)\neq\emptyset$. This relation is reflexive and transitive, but in general is not antisymmetric.
Definition 5.1.
A small skeletal category ${\mathscr{C}}$ is called directed if the relation $\leqslant$ on $\operatorname{Ob}({\mathscr{C}})$ defined above is a partial order. A directed category ${\mathscr{C}}$ is called noetherian if the poset $(\operatorname{Ob}({\mathscr{C}}),\leqslant)$ is noetherian; that is, ${\mathscr{C}}$ does not contain an infinite sequence of morphisms
[…]
such that all objects appearing in it are distinct. Dually, one defines artinian directed categories.
Remark 5.2.
When ${\mathscr{C}}$ is a directed category, one can construct a natural filtration on $\operatorname{Ob}({\mathscr{C}})$ as follows. Explicitly, let $O_{0}=\emptyset$, $O_{1}$ be the set of objects in $\operatorname{Ob}({\mathscr{C}})$ minimal with respect to $\leqslant$. For an ordinal $\chi$, let $O_{\chi}$ be the union of $O_{\chi-1}$ and the set of minimal elements in $\operatorname{Ob}({\mathscr{C}})\backslash O_{\chi-1}$ provided that $\chi$ is a successor ordinal, and let $O_{\chi}$ be the union of $O_{\lambda}$ for $\lambda<\chi$ if $\chi$ is a limit ordinal. In this way we obtain a chain $O_{1}\subseteq O_{2}\subseteq\cdots\subseteq\operatorname{Ob}({\mathscr{C}})$. Note that ${\mathscr{C}}$ is artinian if and only if there exists an ordinal $\zeta$ such that $O_{\zeta}=\operatorname{Ob}({\mathscr{C}})$. Dually, by taking maximal objects one gets another filtration $\{O^{\prime}_{\chi}\}$ on $\operatorname{Ob}({\mathscr{C}})$, and ${\mathscr{C}}$ is noetherian if and only if there exists an ordinal $\zeta$ such that $O^{\prime}_{\zeta}=\operatorname{Ob}({\mathscr{C}})$. For details, see [8, Subsection 3.1].
Definition 5.3.
Let ${\mathscr{C}}$ be a directed category. A full subcategory $\mathscr{D}$ is called a coideal of ${\mathscr{C}}$ if it is downward closed: given $x\in\operatorname{Ob}(\mathscr{D})$ and $y\in\operatorname{Ob}({\mathscr{C}})$, if $y\leqslant x$, then $y\in\operatorname{Ob}(\mathscr{D})$ as well. Dually, one can define ideals of ${\mathscr{C}}$.
The above concepts generalize downward sets and upward sets of posets.
Definition 5.4.
Let ${\mathscr{C}}$ be a directed category. We say that $x\in\operatorname{Ob}({\mathscr{C}})$ is downward unbounded if there exists a descending chain
$x>x_{1}>x_{2}>\ldots$
of infinite length. Otherwise, $x$ is called downward bounded. Upward unbounded objects and upward bounded objects are defined similarly.
Let ${\mathscr{C}}$ be a directed category. Denote by ${\mathscr{C}}_{db}$ (resp., ${\mathscr{C}}_{du}$) the full subcategory of ${\mathscr{C}}$ consisting of downward bounded (resp., downward unbounded) objects. Similarly, one defines ${\mathscr{C}}_{ub}$ and ${\mathscr{C}}_{uu}$. The following result can be easily checked, so we omit its proof. The reader can also formulate a dual result for the upward bounded/unbounded case.
Lemma 5.5.
Let ${\mathscr{C}}$ be a directed category and $x$ an object in ${\mathscr{C}}$. One has:
  1. (1)
    ${\mathscr{C}}$ is artinian if and only if ${\mathscr{C}}={\mathscr{C}}_{db}$;
  2. (2)
    $x$ is downward bounded if and only if ${\mathscr{C}}(y,x)=\emptyset$ for every downward unbounded object $y$;
  3. (3)
    $x$ is downward unbounded if and only if ${\mathscr{C}}(x,y)=\emptyset$ for every downward bounded object $y$;
  4. (4)
    ${\mathscr{C}}_{du}$ (resp., ${\mathscr{C}}_{db}$) is an ideal (resp., a coideal) of ${\mathscr{C}}$.
From this lemma we deduce an immediate result: if $x$ is downward bounded (resp., downward unbounded), then ${\mathscr{C}}_{db}(-,x)={\mathscr{C}}(-,x)$ (resp., ${\mathscr{C}}_{du}(x,-)={\mathscr{C}}(x,-)$) as sets of morphisms.
Now we introduce the rigid topology and the main theorem of this section. Let ${\mathscr{C}}$ be a small skeletal category, and $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Recall from [13, C2.2] that $x\in\operatorname{Ob}({\mathscr{C}})$ is called $J$-irreducible if $J(x)=\{{\mathscr{C}}(x,-)\}$. For $y\in\operatorname{Ob}({\mathscr{C}})$, let $S_{y}$ be the sieve generated by the family of morphisms (in ${\mathscr{C}}$ rather than ${\mathscr{C}^{\mathrm{op}}}$) from $y$ to all $J$-irreducible objects. It can be explicitly described as follows
$S_{y}=\bigcup_{x\text{ irreducible}}{\mathscr{C}}(x,-)\circ{\mathscr{C}}(y,x)$
where ${\mathscr{C}}(x,-)\circ{\mathscr{C}}(y,x)=\{f\circ g\mid f:x\to\bullet\text{ and }g:y\to x\}$.
Definition 5.6.
A Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is called rigid if $S_{y}\in J(y)$ for every $y\in\operatorname{Ob}({\mathscr{C}})$.
A rigid Grothendieck topology $J$ is also called a subcategory topology with respect to the full subcategory consisting of $J$-irreducible objects. When $J$ is a rigid topology, $S_{y}$ is the unique minimal covering sieve in $J(y)$, so $J(y)=\{S\subseteq{\mathscr{C}}(y,-)\mid S\supseteq S_{y}\}$. For details, see [13, C2.2, Example 2.2.4(d), Remark 2.2.5, and Definition 2.2.18].
The main theorem of this section is:
Theorem 5.7.
Let ${\mathscr{C}}$ be a directed category. Then every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if ${\mathscr{C}}$ is a noetherian EI category.
Remark 5.8.
[…]
The proof of the above theorem is complicated and lengthy, so we give it in the next two subsections.
5.2. The necessary direction
In this subsection we prove the only if direction of Theorem 5.7. Throughout this subsection let ${\mathscr{C}}$ be a directed category. As the starting point, we consider the special case that ${\mathscr{C}}$ has only one object, so we can identify ${\mathscr{C}^{\mathrm{op}}}$ with a monoid $M$, and correspondingly, a sieve on ${\mathscr{C}^{\mathrm{op}}}$ is nothing but a right ideal of $M$.
Lemma 5.9.
A monoid $M$ has exactly two Grothendieck topologies if and only if $M$ is a group.
Proof.
A group $M$ only has two right ideals; that is, $M$ itself and the empty set. Consequently, there are exactly two Grothendieck topologies on $M$: the trivial topology and the maximal topology.
Conversely, if $M$ is not a group, let $N\subseteq M$ be the subset consisting of elements $m$ which are not right invertible. It is easy to check that $N$ is a nonempty proper right ideal of $M$. Furthermore, it is contained in the dense topology $J_{d}$ on $M$, which by definition is the set of right ideals $S$ of $M$ satisfying the following condition: for any $m\in M$, there is a certain $n\in M$ such that $mn\in S$. Indeed, given $m\in M$, if $m$ is not right invertible, then $m\cdot 1\in N$; if $m$ is right invertible, by taking a right inverse $n\in M$ of $m$ and a certain $l\in N$, then $m(nl)=l\in N$.
Note that $J_{d}$ is different from the maximal topology since it does not contain the empty set. It is also different from the trivial topology since it contains proper subset $N$ of $M$. Therefore, there are at least three different Grothendieck topologies on $M$. ∎
Now we describe two technical lemmas playing a key role for proving the necessary direction of Theorem 5.7.
Lemma 5.10.
Let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ and ${\mathscr{D}}$ an ideal of ${\mathscr{C}}$. Then the restriction of $J$ on ${\mathscr{D}}^{{\mathrm{op}}}$ given by
$J_{{\mathscr{D}}}(x)=\{S\cap{\mathscr{D}}(x,-)\mid S\in J(x)\}$
for $x\in\operatorname{Ob}({\mathscr{D}})$ defines a Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$. Moreover, every Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$ can be obtained in this way. Consequently, if every Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is rigid, so is every Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$.
Proof.
Since ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$, we have ${\mathscr{C}}(x,y)=\emptyset$ for $x\in\operatorname{Ob}({\mathscr{D}})$ and $y\notin\operatorname{Ob}({\mathscr{D}})$. Consequently, as sets of morphisms, we have $S\cap{\mathscr{D}}(x,-)=S$ for $S\in J(x)$ and $x\in\operatorname{Ob}({\mathscr{D}})$. With this observation, it is easy to check that $J_{{\mathscr{D}}^{{\mathrm{op}}}}$ is indeed a Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$.
Conversely, given a Grothendieck topology $K$ on ${\mathscr{D}}^{{\mathrm{op}}}$, we define a rule $J$ for ${\mathscr{C}^{\mathrm{op}}}$ by setting $J(x)=K(x)$ for $x\in\operatorname{Ob}({\mathscr{D}})$ and $J(y)=\{{\mathscr{C}}(y,-)\}$ for $y\in\operatorname{Ob}({\mathscr{C}})\setminus\operatorname{Ob}({\mathscr{D}})$. It is routine to check that $J$ is a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$, and furthermore, its restriction on ${\mathscr{D}}^{{\mathrm{op}}}$ is precisely $K$.
To establish the last statement, we need to show that the restriction $J_{{\mathscr{D}}^{{\mathrm{op}}}}$ is a rigid Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$ whenever $J$ is a rigid Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. But this is clear. Indeed, since $J$ is rigid, for every $x\in\operatorname{Ob}({\mathscr{D}})$, morphisms in ${\mathscr{C}}$ from $x$ to $J$-irreducible objects generate a covering sieve $S_{x}\in J(x)$. Since ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$, as explained before, $J_{{\mathscr{D}}^{{\mathrm{op}}}}$-irreducible objects are precisely $J$-irreducible objects contained in $\operatorname{Ob}({\mathscr{D}})$, and only these objects contribute to $S_{x}$. Consequently, $S_{x}=S_{x}\cap{\mathscr{D}}(x,-)\in J_{{\mathscr{D}}^{{\mathrm{op}}}}(x)$ is generated by morphisms in ${\mathscr{D}}$ from $x$ to $J_{{\mathscr{D}}^{{\mathrm{op}}}}$-irreducible objects, so $J_{{\mathscr{D}}^{{\mathrm{op}}}}$ is rigid. ∎
The second lemma gives a dual result: if every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid, then so is every Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$ where $\mathscr{E}$ is a coideal of ${\mathscr{C}}$. To prove this result, we need to introduce a slightly more complicated construction.
Let $\mathscr{E}$ be a coideal of ${\mathscr{C}}$, and let ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of those objects not contained in $\operatorname{Ob}(\mathscr{E})$. It is easy to check that ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$. Let $K$ be the maximal topology on ${\mathscr{D}}^{{\mathrm{op}}}$ and $L$ an arbitrary Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$. We define a rule $J$ on ${\mathscr{C}^{\mathrm{op}}}$ as follows:
$J(x)=\begin{cases}K(x),&x\in\operatorname{Ob}({\mathscr{D}});\\ \{S\subseteq{\mathscr{C}}(x,-)\mid S\cap\mathscr{E}(x,-)\in L(x)\},&x\in\operatorname{Ob}(\mathscr{E}).\end{cases}$
We point out that $S\subseteq{\mathscr{C}}(x,-)$ is contained in $J(x)$ if and only if there is a certain $S^{\prime}\in L(x)$ such that $S^{\prime}\subseteq S$. The only if direction is trivial by taking $S^{\prime}=S\cap\mathscr{E}(x,-)$. The if direction is also clear: $S^{\prime}\subseteq S$ implies that $S\cap\mathscr{E}(x,-)\supseteq S^{\prime}$, so $S\cap\mathscr{E}(x,-)$ is contained in $L(x)$.
Lemma 5.11.
The rule $J$ defined above is a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Furthermore, if all Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are rigid, then so is every Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$.
Proof.
The maximal axiom clearly holds. For the stability axiom, let $S\in J(x)$ be a covering sieve, and $f:x\to y$ be a morphism in ${\mathscr{C}}$. If $y$ is an object in ${\mathscr{D}}$, then $f^{\ast}(S)\subseteq{\mathscr{C}}(y,-)$ is contained in $K(y)=J(y)$ since $K$ is the maximal topology on ${\mathscr{D}}^{{\mathrm{op}}}$ and ${\mathscr{D}}(y,-)={\mathscr{C}}(y,-)$ as ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$. If $y$ is contained in $\mathscr{E}$, so is $x$ because $\mathscr{E}$ is a coideal. Since $S\cap\mathscr{E}(x,-)\in L(x)$ by definition of $J$, in this case
$f^{\ast}(S)\cap\mathscr{E}(y,-)=f^{\ast}(S)\cap f^{\ast}(\mathscr{E}(x,-))=f^{\ast}(S\cap\mathscr{E}(x,-))$
is contained in $L(y)$ by the stability axiom on $L$, and hence $f^{\ast}(S)\in J(x)$ as desired.
To check the transitivity condition, take $T\subseteq{\mathscr{C}}(x,-)$, $S\in J(x)$ such that $f^{\ast}(T)\in J(y)$ for each $f:x\to y$ in $S$. If $x$ is contained in $\mathscr{D}$, then clearly $T\in J(x)$. Otherwise, one has
$T\supseteq\bigcup_{f\in S}f^{\ast}(T)\circ f\supseteq\bigcup_{f\in S\cap\mathscr{E}(x,-)}f^{\ast}(T)\circ f\supseteq\bigcup_{f:x\to y\atop f\in S\cap\mathscr{E}(x,-)}(f^{\ast}(T)\cap\mathscr{E}(y,-))\circ f.$
Denote the right side of the above formula by $T^{\prime}$, which is a subfunctor of $\mathscr{E}(x,-)$ since $x\in\operatorname{Ob}(\mathscr{E})$. Note that $S\cap\mathscr{E}(x,-)\in L(x)$ and $f^{\ast}(T)\cap\mathscr{E}(y,-)\in L(y)$ by definition of $J$, and $f^{\ast}(T)\cap\mathscr{E}(y,-)$ is a subset of $f^{\ast}(T^{\prime})$. It follows that $f^{\ast}(T^{\prime})\in L(y)$ for each $f\in S\cap\mathscr{E}(x,-)$. Applying the transitivity axiom for $L$ to $S\cap\mathscr{E}(x,-)\in L(x)$ and $T^{\prime}\subseteq\mathscr{E}(x,-)$, we deduce that $T^{\prime}\in L(x)$. Consequently, $T\in J(x)$ since it contains the $L$-covering sieve $T^{\prime}$.
Now we prove the second statement. Let $L$ be a Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$ and construct a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ as above. Note that for every $x\in\operatorname{Ob}(\mathscr{E})$ one has
$L(x)=\{S\cap\mathscr{E}(x,-)\mid S\in J(x)\}.$
Indeed, given a covering sieve $T\in L(x)$ in $\mathscr{E}^{{\mathrm{op}}}$, it generates a sieve $S\subseteq{\mathscr{C}}(x,-)$ in ${\mathscr{C}^{\mathrm{op}}}$, which clearly belongs to $J(x)$ since $S\supseteq T$. It is easy to check that $S\cap\mathscr{E}(x,-)=T$, so the left side of the above identity is a subset of the right side. The inclusion of the other direction follows from the construction of $J$.
Since $J$ is rigid, for every $x\in\operatorname{Ob}({\mathscr{C}})$, morphisms in ${\mathscr{C}}$ from $x$ to $J$-irreducible objects form a covering sieve $S_{x}\in J(x)$. Note that by our construction of $J$, every object contained in ${\mathscr{D}}$ is not $J$-irreducible since $K$ is the maximal topology on ${\mathscr{D}}$. Therefore, $S_{x}$ is generated by morphisms in ${\mathscr{C}}$ from $x$ to $J$-irreducible objects contained in $\mathscr{E}$, which coincide with $L$-irreducible objects. Consequently, for $x\in\operatorname{Ob}(\mathscr{E})$, the covering sieve $(S_{x}\cap\mathscr{E}(-,x))\in L(x)$ is generated by morphisms in $\mathscr{E}$ from $x$ to $L$-irreducible objects. Therefore, $L$ is rigid. ∎
We are ready to prove the necessary direction of Theorem 5.7.
Proposition 5.12.
Let ${\mathscr{C}}$ be a directed category. If every Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is rigid, then ${\mathscr{C}}$ is a noetherian EI category.
Proof.
We first show that ${\mathscr{C}}$ is an EI category; that is, for every $x\in\operatorname{Ob}({\mathscr{C}})$, the monoid $M_{x}={\mathscr{C}}(x,x)$ is a group. Let ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of objects $y$ with $y\geqslant x$. Clearly, ${\mathscr{D}}$ is an ideal of ${\mathscr{C}}$. By Lemma 5.10, every Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$ shall be rigid. Now let $\mathscr{E}$ be the full subcategory of ${\mathscr{D}}$ consisting of the single object $x$. Then $\mathscr{E}$ is a coideal of ${\mathscr{D}}$. By Lemma 5.11, every Grothendieck topology on $\mathscr{E}^{{\mathrm{op}}}$ shall be rigid as well. By Lemma 5.9, $M_{x}^{{\mathrm{op}}}$ is a group, so is $M_{x}$. Therefore, ${\mathscr{C}}$ is an EI category.
Next we show that ${\mathscr{C}}$ is noetherian by contradiction. If this is not true, then the full subcategory ${\mathscr{C}}_{uu}$ consisting of upward unbounded objects is nonempty, and is a coideal of ${\mathscr{C}}$ by the dual version of Lemma 5.5. By Lemma 5.11, every Grothendieck topology on ${\mathscr{C}}_{uu}^{{\mathrm{op}}}$ shall be rigid. We claim that the dense topology $J_{d}$ on ${\mathscr{C}}_{uu}^{{\mathrm{op}}}$ is not rigid, so the conclusion follows by this contradiction.
Given $x\in\operatorname{Ob}({\mathscr{C}}_{du})$, define
$T_{x}=\bigsqcup_{y\in\operatorname{Ob}({\mathscr{C}}_{uu})\atop y\neq x}{\mathscr{C}}(x,y),$
a proper subfunctor of ${\mathscr{C}}_{uu}(x,-)$. Since $x$ is upward unbounded, there are infinitely many objects $y\in\operatorname{Ob}({\mathscr{C}}_{uu})$ such that ${\mathscr{C}}_{uu}(x,y)$ is nonempty, so $T_{x}$ is nonempty. Moreover, it is easy to see that $T_{x}\in J_{d}(x)$. Therefore, $x$ is not $J_{d}$-irreducible, so no object in ${\mathscr{C}}_{uu}$ is $J_{d}$-irreducible. Consequently, $J_{d}$ is not rigid because the subcategory topology $J$ on ${\mathscr{C}}_{uu}$ such that no object is $J$-irreducible is the maximal topology, which is different from $J_{d}$. ∎
5.3. Sufficient direction
In this subsection we prove the sufficient direction of Theorem 5.7. Throughout this subsection let ${\mathscr{C}}$ be a noetherian EI category unless otherwise specified. We firstly show that for every Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ and every $x\in\operatorname{Ob}({\mathscr{C}})$, $J(x)$ has a unique minimal element: the intersection of all $S\in J(x)$.
Lemma 5.13.
Every Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is closed under arbitrary intersections.
Proof.
Since $J$ is closed under inclusions, it suffices to show
$\bigcap_{S\in J(x)}S\in J(x)$
for all $x\in\operatorname{Ob}({\mathscr{C}})$. If the conclusion is false, then the set $O$ consisting of objects $x$ such that the above formula fails is not empty. Since $\operatorname{Ob}({\mathscr{C}})$ is a noetherian poset, we can find a maximal object $x$ in $O$. We deduce that:
  • $\emptyset\notin J(x)$, since otherwise the formula holds;
  • $x$ is not a maximal object in ${\mathscr{C}}$, since otherwise ${\mathscr{C}}(x,-)$ has only two subfunctors, namely itself and the empty functor, and the formula also holds;
  • the cardinality $|J(x)|$ must be infinity since Grothendieck topologies are closed under finite intersections.
Take an arbitrary $T$ in $J(x)$. We may assume that $T$ is different from ${\mathscr{C}}(x,-)$ (this is possible since $J(x)$ is an infinite set). By the transitivity axiom, there must exist a certain $f:x\to y$ in $T$ such that
$f^{\ast}(\bigcap_{S\in J(x)}S)=\bigcap_{S\in J(x)}f^{\ast}(S)\notin J(y).$
But clearly we have
$\bigcap_{S\in J(x)}f^{\ast}(S)\supseteq\bigcap_{T\in J(y)}T.$
This forces
$\bigcap_{T\in J(y)}T\notin J(y),$
that is, $y\in O$ as well. However, because ${\mathscr{C}}$ is a skeletal EI category, $f$ cannot be an endomorphism since otherwise $f$ is an automorphism and hence $T={\mathscr{C}}(x,-)$. So, we have $x<y$, contradicting the assumption that $x$ is maximal in $O$. The conclusion then follows by contradiction. ∎
By this lemma, a Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ determines a family of minimal covering sieves each of which belongs to $J(x)$ for a distinct $x\in\operatorname{Ob}({\mathscr{C}})$. Now we consider the converse procedure, generalizing the idea in [19] to construct Grothendieck topologies on ${\mathscr{C}}^{{\mathrm{op}}}$ by giving a family of minimal covering sieves satisfying a special requirement. We assign to each object $x$ in ${\mathscr{C}}$ a subfunctor $S_{x}\subseteq{\mathscr{C}}(x,-)$, and call this assignment consistent if the following condition holds: for every $x\in\operatorname{Ob}({\mathscr{C}})$, either $S_{x}={\mathscr{C}}(x,-)$ or
$S_{x}=\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})\atop f\in{\mathscr{C}}(x,y)}S_{y}f=\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})}S_{y}\circ{\mathscr{C}}(x,y)$ (5.1)
where $S_{y}f=\{gf\mid g\in S_{y}\}$. By convention, the above union is empty when the index is empty. Such a consistent family $\{S_{x}\}_{x\in\operatorname{Ob}({\mathscr{C}})}$ determines a rule $J$ by
$J(x)=\{S\subseteq{\mathscr{C}}(x,-)\mid S\supseteq S_{x}\}.$
Proposition 5.14.
The rule $J$ defined above is a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$, and this construction exhausts all Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$.
Proof.
We show that $J$ is indeed a Grothendieck topology on ${\mathscr{C}}^{{\mathrm{op}}}$. The maximal axiom holds obviously. For the stability axiom, we observe that if $S\in J(x)$, then $S\supseteq S_{x}$, and for each morphism $f:x\to y$ in ${\mathscr{C}}$, one has
$f^{\ast}(S)\supseteq f^{\ast}(S_{x})\supseteq f^{\ast}(S_{y}f)\supseteq S_{y},$
so $f^{\ast}(S)\in J(y)$. It remains to check the transitivity axiom.
Take a subfunctor $S\subseteq{\mathscr{C}}(x,-)$ and $T\in J(x)$ such that $f^{\ast}(S)\in J(y)$ for each morphism $f:x\to y$ in $T$. We want to check that $S\in J(x)$; that is, $S_{x}\subseteq S$. If $S={\mathscr{C}}(x,-)$, the conclusion holds trivially. Similarly, if $T$ contains an endomorphism on $x$, then $T={\mathscr{C}}(x,-)$, so by taking $f$ to be the identity morphism on $x$ we have $S=f^{\ast}(S)\in J(x)$. Now suppose that both $S$ and $T$ do not contain any endomorphism on $x$. In this case it suffices to show the following claim:
(§)§($\S$)( § ) each $\alpha\in S_{x}$ can be written as $\alpha=\beta\gamma$ with $\beta\in S_{y}$ and $\gamma\in T$.
Indeed, since for each $\gamma\in T$ one has $\gamma^{\ast}(S)\in J(y)$, namely $\gamma^{\ast}(S)\supseteq S_{y}$, we deduce that $S\supseteq S_{y}T$. Thus if the above claim is valid, then $\alpha\in S_{y}T\subseteq S$ for each $\alpha\in S_{x}$, so $S_{x}\subseteq S$ as desired.
Now we check the claim (§)§($\S$)( § ). Since $S\neq{\mathscr{C}}(x,-)$, by the consistent condition we can write $\alpha$ as a composite $\beta_{1}\gamma_{1}$ with $\gamma_{1}\in{\mathscr{C}}(x,x_{1})$ for a certain object $x_{1}\neq x$ and $\beta_{1}\in S_{x_{1}}$. We have two cases:
  • If $S_{x_{1}}={\mathscr{C}}(x_{1},-)$, then $\gamma_{1}\in S_{x}\subseteq T$ by the consistent condition, so the claim holds.
  • Otherwise, equation (5.1) holds for $S_{x_{1}}$. Thus one can replace $x$ by $x_{1}$ and write $\beta_{1}=\beta_{2}\gamma_{2}$ where $\gamma_{2}\in{\mathscr{C}}(x_{1},x_{2})$ and $\beta_{2}\in S_{x_{2}}$. Clearly, if $S_{x_{2}}={\mathscr{C}}(x_{2},-)$, then applying the argument as in the previous case, we deduce the claim.
Continuing the above procedure we obtain a sequence of objects $x<x_{1}<x_{2}<\ldots$ which must be finite since ${\mathscr{C}}$ is noetherian. Let $x_{n}$ be the last object, then either $S_{x_{n}}={\mathscr{C}}(x_{n},-)$ (and hence (§)§($\S$)( § ) holds) or $x_{n}$ is maximal with respect to $\leqslant$. For the second case, we have $\beta_{n}\in S_{x_{n}}$ is an automorphism since ${\mathscr{C}}(x_{n},x_{n})$ is assumed to be a group, so $S_{x_{n}}={\mathscr{C}}(x_{n},-)$ as well. Consequently, the claim (§)§($\S$)( § ) holds in both cases.
Now we check the second statement. Let $J$ be an arbitrary Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$, and for every $x\in\operatorname{Ob}({\mathscr{C}})$ let $S_{x}$ be the intersection of all members in $J(x)$, which is contained in $J(x)$ by Lemma 5.13. It remains to show that the family $\{S_{x}\mid x\in\operatorname{Ob}({\mathscr{C}})\}$ is consistent; that is, equation (5.1) holds provided that $S_{x}\neq{\mathscr{C}}(x,-)$. When $x$ is a maximal object, ${\mathscr{C}}(x,-)$ has only two subfunctors, namely itself and the empty functor. In this case it is easy to check the consistent condition.
Suppose that $x$ is not a maximal object. For each morphism $f:x\to y$ with $x\neq y$, one has $f^{\ast}(S_{x})\in J(y)$, so $f^{\ast}(S_{x})\supseteq S_{y}$, namely $S_{x}\supseteq S_{y}f$. Consequently, we obtain
$S_{x}\supseteq\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})\atop f\in{\mathscr{C}}(x,y)}S_{y}f.$
To check the inclusion of the other direction, it suffices to show that the right side, denoted by $T$, is contained in $J(x)$. We prove it by the transitivity axiom. Indeed, for an arbitrary $g:x\to z$ in $S_{x}$, we have $z\neq x$ since $S_{x}\neq{\mathscr{C}}(x,-)$. Thus $S_{z}g\subseteq T$ and hence
$g^{\ast}(T)\supseteq g^{\ast}(S_{z}g)\supseteq S_{z}\in J(z).$
Since $J$ is closed under inclusions, we deduce that $g^{\ast}(T)\in J(z)$ as well. The transitivity axiom then tells us that $T\in J(x)$, as desired. ∎
The reader may guess that the Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ constructed before Proposition 5.14 is rigid, or equivalently, each $S_{x}$ in the consistent family is generated by morphisms in ${\mathscr{C}}$ from $x$ to $J$-irreducible objects. This is indeed the case by the following corollary, which establishes the sufficient direction of Theorem 5.7.
Corollary 5.15.
Every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is rigid.
Proof.
Proposition 5.14 tells us that every Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is determined by a consistent family. By comparing formula (5.1) and Definition 5.6, it suffices to show that each $S_{x}$ in formula (5.1) coincides with the subfunctor $S^{\prime}_{x}\subseteq{\mathscr{C}}(x,-)$ generated by all morphisms from $x$ to $J$-irreducible objects. When $S_{x}={\mathscr{C}}(x,-)$, $x$ is $J$-irreducible, so this is trivially true.
Suppose that $x$ is not $J$-irreducible. In this case we have
$S_{x}=\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})\atop f\in{\mathscr{C}}(x,y)}S_{y}f=\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})}S_{y}\circ{\mathscr{C}}(x,y)$
and
$S_{x}^{\prime}=\bigcup_{y\text{ irred.}}{\mathscr{C}}(y,-)\circ{\mathscr{C}}(x,y).$
Noth that $S_{y}={\mathscr{C}}(y,-)$ when $y$ is $J$-irreducible, so $S_{x}^{\prime}\subseteq S_{x}$. We prove the other inclusion $S_{x}\subseteq S_{x}^{\prime}$ via a transfinite induction based on a filtration $\{O^{\prime}_{\chi}\}$ of maximal objects specified in Remark 5.2. Explicitly, there is an increasing sequence $O_{1}^{\prime}\subseteq O_{2}^{\prime}\subseteq O_{3}^{\prime}\subseteq\ldots$ of subsets of $\operatorname{Ob}({\mathscr{C}})$ satisfying the following conditions:
  • there exists an ordinal $\zeta$ such that $\operatorname{Ob}({\mathscr{C}})=O_{\zeta}^{\prime}$ since ${\mathscr{C}}$ is noetherian;
  • ${\mathscr{C}}(z,w)=\emptyset$ if $w\in O^{\prime}_{\lambda+1}\setminus O_{\lambda}^{\prime}$ and $z\in O^{\prime}_{\lambda}$ for a certain ordinal $\lambda$;
  • for $x\in\operatorname{Ob}({\mathscr{C}})$, there is a minimal ordinal $d(x)$ such that $x\in O^{\prime}_{d(x)}$.
If $x$ is maximal with respect to the partial order $\leqslant$, i.e., $d(x)=1$, then
$S_{x}=\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})\atop f\in{\mathscr{C}}(x,y)}S_{y}f=\emptyset$
since there is no morphism $f:x\to y$ with $x\neq y$. Clearly, $S_{x}\subseteq S_{x}^{\prime}$.
Suppose that $d(x)=\lambda$. Since $x$ is not $J$-irreducible, one has
$S_{x}=(\bigcup_{x\neq y\atop y\text{ irred.}}S_{y}\circ{\mathscr{C}}(x,y))\cup(\bigcup_{x\neq y\atop y\text{ not irred.}}S_{y}\circ{\mathscr{C}}(x,y))\subseteq S_{x}^{\prime}\cup(\bigcup_{x\neq y\atop y\text{ not irred.}}S_{y}\circ{\mathscr{C}}(x,y)).$
But $x\neq y$ and ${\mathscr{C}}(x,y)\neq\emptyset$ imply $d(y)<d(x)$. By the induction hypothesis, $S_{y}\subseteq S_{y}^{\prime}$, and hence
$S_{y}\circ{\mathscr{C}}(x,y)\subseteq S_{y}^{\prime}\circ{\mathscr{C}}(x,y)=\bigcup_{z\text{ irred.}}{\mathscr{C}}(z,-)\circ{\mathscr{C}}(y,z)\circ{\mathscr{C}}(x,y)\subseteq\bigcup_{z\text{ irred.}}{\mathscr{C}}(z,-)\circ{\mathscr{C}}(x,z)\subseteq S_{x}^{\prime}$
when $y$ is not $J$-irreducible. Consequently, one has $S_{x}\subseteq S_{x}^{\prime}$. The conclusion then follows from a transfinite induction. ∎
When ${\mathscr{C}}$ is a noetherian EI category, the following result gives a very transparent description for $J$-torsion $\mathcal{O}$-modules and categories of sheaves of modules.
Theorem 5.16.
Let ${\mathscr{C}}$ be a noetherian EI category, $J$ a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$, $\mathcal{O}$ a structure sheaf on $\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}}$, and $\mathscr{D}$ the full subcategory of ${\mathscr{C}}$ consisting of $J$-irreducible objects. Then:
  1. (1)
    an $\mathcal{O}$-module $V$ lies in $\mathcal{T}(J)$ if and only if $V_{x}=0$ for each $x\in\operatorname{Ob}({\mathscr{D}})$;
  2. (2)
    consequently, one has
    ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})\simeq\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}$
    where $\mathcal{O}_{{\mathscr{D}}}$ is the structure presheaf on ${\mathscr{D}}^{{\mathrm{op}}}$ obtained via restricting $\mathcal{O}$ to ${\mathscr{D}}$.
Proof.
(1): The only if direction is clear. For the other direction, take an $\mathcal{O}$-module $V$ such that $V_{x}=0$ for every $x\in\operatorname{Ob}({\mathscr{D}})$, an arbitrary $x\in\operatorname{Ob}({\mathscr{C}})$ and an element $v\in V_{x}$, we want to show that $S_{x}\cdot v=0$. If $S_{x}={\mathscr{C}}(x,-)$, then $V_{x}=0$ and hence $v=0$. If $S_{x}=\emptyset$, then this trivially holds by our convention, so we may assume that $S_{x}\neq\emptyset$. By the consistent condition,
$S_{x}=\bigcup_{x\neq y\in\operatorname{Ob}({\mathscr{C}})\atop f\in{\mathscr{C}}(x,y)}S_{y}f.$
Take a morphism $(g:x\to z)\in S_{x}$ and write it as $g=g_{1}f$ with $f:x\to x_{1}$ and $g_{1}:x_{1}\to z$ in $S_{x_{1}}$. Then $g\cdot v=g_{1}\cdot(f\cdot v)$. Replace $v$ by $f\cdot v$, $x$ by $x_{1}$, $g$ by $g_{1}$ and repeat the above argument. This procedure ends with two cases:
  • We obtain a certain $x_{n}\neq z$ such that $S_{x_{n}}={\mathscr{C}}(x_{n},-)$. In this case the conclusion holds since $V_{x_{n}}=0$.
  • After finitely many steps we obtain $x_{n}=z$ since ${\mathscr{C}}$ is noetherian. In this case, $g_{n}$ is a morphism from $z$ to itself, and hence an automorphism. Since $S_{z}$ contains $g_{n}$, it coincides with ${\mathscr{C}}(z,-)$, and the conclusion also holds since $V_{z}=0$.
This finishes the proof of the first statement.
(2): This is essentially [13, C2.2, Theorem 2.2.3]. For the convenience of the reader, we give a sketched proof. Given an $\mathcal{O}$-module $V$, its restriction to ${\mathscr{D}}$ is clearly an $\mathcal{O}_{{\mathscr{D}}}$-module. In this way we obtain an restriction functor $\textsl{res}:\mathcal{O}\operatorname{-Mod}\to\mathcal{O}_{\mathscr{D}}\operatorname{-Mod}$. It has a right adjoint given by the right Kan extension described as follows: given an $\mathcal{O}_{{\mathscr{D}}}$-module $W$ and an object $x\in\operatorname{Ob}({\mathscr{C}})$,
[…]
where $\iota:{\mathscr{D}}\to{\mathscr{C}}$ is the inclusion functor, and the limit is taken over the comma category $(\iota\downarrow\mathrm{const}_{x})$. When $x\in\operatorname{Ob}({\mathscr{D}})$, it is not hard to see that $(\iota\downarrow\mathrm{const}_{x})$ is a coslice category whose objects are morphisms in ${\mathscr{D}}$ ending at $x$. In this situation, the identity morphism on $x$ is the initial object of the coslice category. It follows that $(\mathrm{Ran}_{\iota}W)_{x}\cong W_{x}$ for $x\in\operatorname{Ob}({\mathscr{D}})$. Consequently, $\textsl{res}\circ\mathrm{Ran}_{\iota}$ is isomorphic to the identity functor on $\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}$. By [9, Proposition III.2.5], we conclude that
$\mathcal{O}_{{\mathscr{D}}}\operatorname{-Mod}\simeq\mathcal{O}\operatorname{-Mod}/\ker(\textsl{res})=\mathcal{O}\operatorname{-Mod}/\mathcal{T}(J)\simeq{\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\,\mathcal{O})$
since the kernel of res is precisely $\mathcal{T}(J)$ by (1). ∎
Remark 5.17.
Note that the above result holds for any rigid Grothendieck topology $J$ on any small category ${\mathscr{C}}^{{\mathrm{op}}}$ by letting ${\mathscr{D}}$ be the full subcategory of ${\mathscr{C}}$ consisting of $J$-irreducible objects; see [13, C2.2, Theorem 2.2.3] and [13, C2.2, Example 2.2.4(d)].
6. Grothendieck topologies on certain noetherian EI categories
We have classified all Grothendieck topologies on ${\mathscr{C}}^{{\mathrm{op}}}$ when ${\mathscr{C}}$ is a noetherian EI category. The question becomes much more complicated for arbitrary EI categories, even under the assumption that ${\mathscr{C}}$ is an artinian EI categories. Throughout this section let ${\mathscr{C}}$ be an EI category satisfying the following conditions:
  • objects in ${\mathscr{C}}$ are parameterized by ${\mathbb{N}}$, so by abuse of notation we denote them by $n$, $n\in{\mathbb{N}}$;
  • ${\mathscr{C}}(m,n)\neq\emptyset$ if and only if $m\leqslant n$;
  • ${\mathscr{C}}(n,n)$ acts transitively on ${\mathscr{C}}(m,n)$ for every pair of objects $m$ and $n$;
  • ${\mathscr{C}}$ is a graded category equipped with a rank function $\deg:\mathrm{Mor}({\mathscr{C}})\to{\mathbb{N}}$ such that $\deg(f)=n-m$ for every morphism $f:m\to n$;
  • $\mathrm{Mor}({\mathscr{C}})$ is generated in degrees 0 and 1, namely every $f:m\to n$ with $n>m$ can be written as a composite of $n-m$ morphisms of degree 1.
We call them EI categories of type ${\mathbb{N}}$. Typical examples include the following categories:
  • the posets ${\mathbb{N}}$,
  • skeletal full subcategories of the category ${\mathrm{FI}}$ of finite sets and injections,
  • skeletal full subcategories of the category $\mathrm{VI}(\mathbbm{k})$ of finite dimensional vector spaces over a field $\mathbbm{k}$ and linear injections,
  • skeletal full subcategories of the category of finite cyclic $p$-groups and injective homomorphisms, where $p$ is a prime.
It is clear that ${\mathscr{C}}$ is an artinian EI category. Our goal is to classify Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$.
Lemma 6.1.
For $m\in{\mathbb{N}}$, every nonempty subfunctor $S\subseteq{\mathscr{C}}(m,-)$ is principal; that is, $S$ is generated by a single morphism $f:m\to n$ in ${\mathscr{C}}$.
Proof.
Choose a morphism $f:m\to n$ in $S$ such that $\deg(f)$ is minimal. We claim that
$S=\{gf\mid g\in{\mathscr{C}}(n,-)\}.$
Clearly, the right side is a subset of $S$. On the other hand, for every $h:m\to r$ in $S$, one has $\deg(h)\geqslant\deg(f)$, so $r\geqslant n$. Therefore, one can decompose $h$ as follows:
$m\overset{h_{1}}{\longrightarrow}m+1\overset{h_{2}}{\longrightarrow}\ldots\overset{h_{n-m}}{\longrightarrow}n\overset{h_{n-m+1}}{\longrightarrow}\ldots\overset{h_{n-r}}{\longrightarrow}r$
Since ${\mathscr{C}}(n,n)$ acts transitively on ${\mathscr{C}}(m,n)$, we can find a certain $\delta\in{\mathscr{C}}(n,n)$ such that
$h=(h_{n-r}\ldots h_{n-m+1})\circ(h_{n-m}\ldots h_{1})=(h_{n-r}\ldots h_{n-m+1})\circ(\delta\circ f)$
contained in $\{gf\mid g\in{\mathscr{C}}(n,-)\}$, so the equality holds. ∎
By the transitive action condition, if $S\subseteq{\mathscr{C}}(n,-)$ contains a morphism of degree $d$, then it contains all morphisms from $n$ to $n+d$. This observation as well as the above lemma tells us that nonempty subfunctors of ${\mathscr{C}}(n,-)$ are parameterized by natural numbers. Explicitly, for $r\in{\mathbb{N}}$, there is a unique subfunctor $S(n,r)\subseteq{\mathscr{C}}(n,-)$ consisting of all morphisms $f:n\to\bullet$ with $\deg(f)\geqslant r$. Moreover, one has the following inclusions:
${\mathscr{C}}(n,-)=S(n,0)\supsetneqq S(n,1)\supsetneqq S(n,2)\supsetneqq\ldots.$
Let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ such that each $J(n)$ does not contain the empty subfunctor. We call it a generic Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Define a function
$\boldsymbol{d}:\mathbb{N}\to\mathbb{N}\cup\{\infty\},\quad n\mapsto d_{n}=\sup\{r\in{\mathbb{N}}\mid S(n,r)\in J(n)\}.$
Note that $d_{n}$ might be $\infty$. Moreover, $J$ is completely determined by $\boldsymbol{d}$; that is,
$J(n)=\{S(n,r)\mid r\leqslant d_{n}\}$
since $J$ is closed under inclusions. Conversely, given such a function $\boldsymbol{d}$, we get a rule $J$ assigning $J(n)$ to $n\in\operatorname{Ob}({\mathscr{C}})$, though $J$ might not be a Grothendieck topology.
The above construction gives an injective map
{generic Grothendieck topology J on $\{\text{generic Grothendieck topology $J$ on }{\mathscr{C}^{\mathrm{op}}}\}\longrightarrow\{\text{function }\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})={\mathbb{N}}\to{\mathbb{N}}\cup\{\infty\}\}.$
Lemma 6.2.
Let $f:m\to n$ be a morphism in ${\mathscr{C}}$, Then one has
$f^{\ast}(S(m,r))=\begin{cases}S(n,r-\deg(f))=S(n,r+m-n),&\text{if }r+m\geqslant n;\\ {\mathscr{C}}(n,-),&\text{else.}\end{cases}$
Proof.
This can be easily verified by the definition. ∎
The following result classifies all generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$.
Proposition 6.3.
Generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by functions
$\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})={\mathbb{N}}\to{\mathbb{N}}\cup\{\infty\}$
satisfying the following condition: if $d_{n}\neq 0$, then $d_{n+1}=d_{n}-1$.
More explicitly, a sequence $\boldsymbol{d}:{\mathbb{N}}\to{\mathbb{N}}\cup\{\infty\}$ satisfies the above condition if and only if it decomposes into the following pieces:
  1. (1)
    $(0,0,0,\ldots,0)$;
  2. (2)
    $(r,r-1,r-2,\ldots,1,0)$;
  3. (3)
    (,,,)($\infty$,$\infty$,$\infty$,$\ldots$)( ∞ , ∞ , ∞ , … ),
where the last piece might or might not appear. But if it appears, it must be the tail of the sequence.
For instances, if all pieces of $\boldsymbol{d}$ are of the first from, then it corresponds to the minimal topology; if $\boldsymbol{d}$ has only one piece which is of the third form, then it corresponds to the atomic Grothendieck topology.
Proof.
First we check that the rule $J$ determined by a function $\boldsymbol{d}$ satisfying the specified condition is indeed a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ (it is clearly generic by the construction). The maximal axiom is clear since ${\mathscr{C}}(n,-)\in J(n)$, so we check the stability axiom and the transitivity axiom.
Stability. Take $m\in\operatorname{Ob}({\mathscr{C}})$, $S(m,r)\in J(m)$ with $r\leqslant d_{m}$, and $f\in{\mathscr{C}}(m,n)$. If $d_{n}=\infty$, then $f^{\ast}(S(m,r))\in J(n)$ since $J(n)$ contains all nonempty subfunctors of ${\mathscr{C}}(n,-)$. Thus we may assume $d_{n}\neq\infty$ and hence $d_{m}\neq\infty$ as well. We have two cases:
  • If $r+m\geqslant n$, then $d_{m}+m\geqslant n$, so $d_{n}=d_{m}+m-n\geqslant 0$. By the previous lemma,
    $f^{\ast}(S(m,r))=S(n,m+r-n)\in J(n)$
    since $m+r-n\leqslant d_{m}+m-n=d_{n}$.
  • If $r+m<n$, then $f\in S(m,r)$ and $f^{\ast}(S(m,r))={\mathscr{C}}(n,-)$, which is clearly contained in $J(n)$.
Transitivity. Take $m\in\operatorname{Ob}({\mathscr{C}})$, $S(m,s)\in J(m)$ with $s\leqslant d_{m}$, and $S(m,r)\subseteq{\mathscr{C}}(m,-)$. Suppose that $f^{\ast}(S(m,r))\in J(n)$ for $f:m\to n$ in $S(m,s)$. We want to show $S(m,r)\in J(m)$; in other words, $r\leqslant d_{m}$. This is clearly true if $d_{m}=\infty$, so we assume that $d_{m}$ is finite. We have two cases:
  1. (1)
    If $d_{m}=0$, then $J(m)=\{{\mathscr{C}}(m,-)\}$ and $S(m,s)={\mathscr{C}}(m,-)$. In this case, one can take $f$ to be the identity morphism on $m$. Therefore,
    $S(m,r)=f^{\ast}(S(m,r))\in J(m)$
    implies $S(m,r)={\mathscr{C}}(m,-)$, so $r=0$ and the conclusion holds.
  2. (2)
    Suppose that $d_{m}>0$. In this case, if $r\leqslant s$, then $r\leqslant d_{m}$, so we can assume that $s<r$. For a morphism $f:m\to m+s$ in $S(m,s)$, by the previous lemma,
    $f^{\ast}(S(m,r))=S(m+s,r-s)\in J(m+s),$
    so $r-s\leqslant d_{m+s}$. But since $d_{m}\geqslant s$, one knows that $d_{m+s}=d_{m}-s$ by the specified condition on $\boldsymbol{d}$. Therefore, we have $r-s\leqslant d_{m}-s$, so $r\leqslant d_{m}$, as desired.
We have checked that the rule $J$ is indeed a generic Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Conversely, given a generic Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$, we want to show that the corresponding function $\boldsymbol{d}$ satisfies the specified condition. One has two cases:
  1. (1)
    If $d_{m}=\infty$, then $S(m,r)\in J(m)$ for $r\geqslant 0$. For $f\in{\mathscr{C}}(m,m+1)$ and $r\geqslant 1$, one has
    $f^{\ast}(S(m,r))=S(m+1,r-1)\in J(m+1),$
    so $J(m+1)$ contains all nonempty subfunctors of ${\mathscr{C}}(m+1,-)$, namely $d_{m+1}=\infty=d_{m}-1$.
  2. (2)
    If $d_{m}<\infty$, one has $d_{m+1}\geqslant d_{m}-1$ by an argument similar to the previous case. If $d_{m+1}>d_{m}-1$, then for $n=m+d_{m}$,
    $d_{n}\geqslant d_{n-1}-1\geqslant\ldots\geqslant d_{m+1}-(n-m-1)\geqslant d_{m}+m+1-n=1,$
    so $S(n,1)\in J(n)$. Consequently, for $f:m\to l$ in $S(m,d_{m})\in J(m)$, one has
    $f^{\ast}(S(m,d_{m}+1))=\begin{cases}S(n,d_{m}+1-\deg(f))=S(n,1)\in J(n),&\text{ if }l=n;\\ S(l,0)\in J(l),&\text{ if }l>n.\end{cases}$
    Therefore, $S(m,d_{m}+1)\in J(m)$ by the transitivity axiom, so $d_{m}\geqslant d_{m}+1$, which is impossible. Thus we also have $d_{m+1}=d_{m}-1$.
Corollary 6.4.
Generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by subsets of $\operatorname{Ob}({\mathscr{C}})$.
Proof.
[…]
$d_{n}=\sup\{s+1\mid f(n)=f(n+1)=\ldots=f(n+s)=1\}.$
Conversely, given a function $\boldsymbol{d}$, we define $f$ by setting $f(n)=1$ when $d_{n}>0$, and $f(n)=0$ when $d_{n}=0$. It is obvious that this construction gives a bijection between $2^{\mathbb{N}}$ and the set of functions $\boldsymbol{d}$ in the above proposition. ∎
Remark 6.5.
Generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$, although are parameterized by subsets of $\operatorname{Ob}({\mathscr{C}})$, need not be rigid. Indeed, the atomic Grothendieck topology $J_{a}$ is generic. But since no object is $J_{a}$-irreducible, it is not rigid. On the other hand, the subcategory topology $K$ on ${\mathscr{C}^{\mathrm{op}}}$ such that no object is $K$-irreducible is the maximal topology, but it is not generic. We also observe that in contrast to Lemma 5.13, the atomic Grothendieck topology $J$ is not closed under arbitrary intersections since for $x\in\operatorname{Ob}({\mathscr{C}})$, the intersection of all $S\in J(x)$ is the empty set not contained in $J(x)$.
Now we consider non-generic Grothendieck topologies $J$ on ${\mathscr{C}^{\mathrm{op}}}$. In this case there exists a certain object $m$ such that $\emptyset\in J(m)$. Here are some elementary observations:
  • By the stability axiom, $\emptyset\in J(n)$ for all $n\geqslant m$.
  • If $m\geqslant 1$ and $\emptyset\notin J(m-1)$, then $J(m-1)$ only contains the subfunctor ${\mathscr{C}}(m-1,-)$. Indeed, if $S(m-1,r)\in J(m-1)$ for a certain $r\geqslant 1$, then $\emptyset\in J(m-1)$ via applying the transitivity axiom to $S(m-1,r)\in J(m-1)$ and $\emptyset\subseteq{\mathscr{C}}(m-1,-)$.
  • There are no objects $n$ such that $J(n)$ consists of all nonempty subfunctors of ${\mathscr{C}}(n,-)$. Indeed, by the first observation, we shall has $n<m$. But applying the transitivity axiom to $\emptyset\subseteq{\mathscr{C}}(n,-)$ and the subfunctor $S(n,m-n)\in J(n)$, we deduce that $\emptyset\in J(n)$, which is impossible.
These observations allow us to define a function
[…]
in the following way: if $J(n)$ contains the empty functor, then $d_{n}=-\infty$; otherwise, $J(n)$ is a finite set by the third observation, so we define $d_{n}$ as in the generic case.
The following result classifies all non-generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$.
Proposition 6.6.
Non-generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by functions
$\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})={\mathbb{N}}\to{\mathbb{N}}\cup\{-\infty\}$
satisfying the following conditions: $d_{n}=-\infty$ for $n\gg 0$, and if $d_{n}\neq 0$, then $d_{n+1}=d_{n}-1$.
More explicitly, a sequence $\boldsymbol{d}:{\mathbb{N}}\to{\mathbb{N}}\cup\{\infty\}$ satisfies the above condition if and only if it decomposes into the following pieces:
  1. (1)
    $(0,0,0,\ldots,0)$;
  2. (2)
    $(r,r-1,r-2,\ldots,1,0)$;
  3. (3)
    (,,,)($\infty$,$\infty$,$\infty$,$\ldots$)( - ∞ , - ∞ , - ∞ , … ),
where the last piece must appear and must be the tail of the sequence.
Proof.
The proof is similar to that of Proposition 6.3, so we omit it. ∎
Corollary 6.7.
Non-generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by finite subsets of $\operatorname{Ob}({\mathscr{C}})$.
Proof.
Note that there is a natural bijection between the set of finite subsets of $\mathbb{N}$ and the set of functions $\{f:\mathbb{N}\to\{0,\,1\}\mid f(n)=0\text{ for }n\gg 0\}$. With this observation, the proof is similar to that of Corollary 6.4. ∎
The following result classifies all rigid Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$.
Proposition 6.8.
A Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ is rigid if and only if its corresponded function $\boldsymbol{d}$ satisfies the following condition: $d_{n}\neq\infty$ for all $n\in{\mathbb{N}}$. In particular, every non-generic Grothendieck topology is rigid.
In other words, the function $\boldsymbol{d}$ can be represented by a sequence which decomposes into the following pieces:
  1. (1)
    $(0,0,0,\ldots,0)$;
  2. (2)
    $(r,r-1,r-2,\ldots,1,0)$;
  3. (3)
    (,,,)($\infty$,$\infty$,$\infty$,$\ldots$)( - ∞ , - ∞ , - ∞ , … ),
where the last piece may or may not appear. But if it appears, it must be the tail of the sequence.
Proof.
It is routine to check that Grothendieck topologies satisfying this condition are rigid. To prove the only if direction, we need to show that if $d_{n}=\infty$ for a certain $n\in{\mathbb{N}}$, then the corresponded Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ is not rigid. Let $n$ be the minimal number such that $d_{n}=\infty$. Note that if $m\geqslant n$, then $d_{m}=\infty$, and hence $m$ is not $J$-irreducible. Thus the subfunctor of ${\mathscr{C}}(n,-)$ generated by morphisms from $n$ to $J$-irreducible objects is the empty set, which is not contained in $J(n)$. Consequently, $J$ is not rigid. ∎
We call a Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ exceptional if it is not rigid, which shall be generic by the previous proposition. Therefore, its corresponds function $\boldsymbol{d}$ satisfies the following condition: $d_{n}=\infty$ for some $n\in{\mathbb{N}}$. Since an explicit description of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ for rigid Grothendieck topologies has been obtained, in the rest of this section we give a description of ${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})$ for exceptional Grothendieck topologies.
Let $J$ be an exceptional Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$ corresponded to a function $\boldsymbol{d}$, and let ${\mathscr{D}}$ be the full subcategory consisting of objects $n$ such that $d_{n}=0$ or $d_{n}=\infty$. Then ${\mathscr{D}}^{{\mathrm{op}}}$ is a $J$-dense subcategory of ${\mathscr{C}}^{{\mathrm{op}}}$ (for a definition of $J$-dense subcategories, see [13, C2.2, Definition 2.2.1]). Therefore, applying the argument in the proof of Theorem 5.16, one has
${\mathrm{Sh}}(\boldsymbol{{\mathscr{C}}}^{{\mathrm{op}}},\mathcal{O})\simeq{\mathrm{Sh}}(\boldsymbol{\mathscr{D}}^{{\mathrm{op}}},\mathcal{O}_{{\mathscr{D}}}),$
where $\boldsymbol{\mathscr{D}}^{{\mathrm{op}}}=({\mathscr{D}}^{{\mathrm{op}}},J_{{\mathscr{D}}})$, $J_{{\mathscr{D}}}$ is a Grothendieck topology on ${\mathscr{D}}^{{\mathrm{op}}}$ such that
$J_{{\mathscr{D}}}(n)=\begin{cases}\{{\mathscr{D}}(n,-)\},&\text{if }d_{n}=0;\\ \{S\mid S\subseteq{\mathscr{D}}(n,-)={\mathscr{C}}(n,-)\},&\text{if }d_{n}=\infty,\end{cases}$
and $\mathcal{O}_{{\mathscr{D}}}$ is the restriction of $\mathcal{O}$ to ${\mathscr{D}}$. The full subcategory ${\mathscr{D}}$ is an artinian category whose objects can still be parameterized by $\mathbb{N}$, and $J_{{\mathscr{D}}}$ corresponds to the sequence $(0,0,\ldots,0,\infty,\infty,\ldots)$.
Let $\mathscr{E}$ be the full subcategory of ${\mathscr{D}}$ consisting of objects $n$ such that $d_{n}=\infty$. It is an ideal of ${\mathscr{D}}$, so the restriction of $J_{{\mathscr{D}}}$ to $\mathscr{E}^{{\mathrm{op}}}$ defines a Grothendieck topology $J_{\mathscr{E}}$ on $\mathscr{E}^{{\mathrm{op}}}$ by Lemma 5.10, and the restriction $\mathcal{O}_{\mathscr{E}}$ of $\mathcal{O}_{{\mathscr{D}}}$ to $\mathscr{E}$ is a structure sheaf on $(\mathscr{E}^{{\mathrm{op}}},J_{\mathscr{E}})$. Note that $J_{\mathscr{E}}$ is the atomic Grothendieck topology.
Proposition 6.9.
[…]
Proof.
This can be checked directly by Definition 2.4. We give a representation theoretic proof. By Corollary 4.3, $V$ is a sheaf of modules on $(\boldsymbol{\mathscr{D}}^{{\mathrm{op}}},\mathcal{O}_{{\mathscr{D}}})$ if and only if the inclusion $\mathbb{Z}S\to\mathbb{Z}{\mathscr{D}}(x,-)$ induces an isomorphism
$\operatorname{Hom}_{\underline{\mathbb{Z}}_{{\mathscr{D}}}\operatorname{-Mod}}(\mathbb{Z}{\mathscr{D}}(x,-),V)\cong\operatorname{Hom}_{\underline{\mathbb{Z}}_{{\mathscr{D}}}\operatorname{-Mod}}(\mathbb{Z}S,V),$ (6.1)
for each $x\in\operatorname{Ob}({\mathscr{D}})$ and each $S\in J_{{\mathscr{D}}}(x)$. If $x\in\operatorname{Ob}({\mathscr{D}})\setminus\operatorname{Ob}(\mathscr{E})$, then $S={\mathscr{D}}(x,-)$, so the above isomorphism automatically holds. When $x\in\operatorname{Ob}(\mathscr{E})$, since $\mathscr{E}$ is an ideal of ${\mathscr{D}}$, we deduce that $J_{{\mathscr{D}}}(x)=J_{\mathscr{E}}(x)$, ${\mathscr{D}}(x,-)=\mathscr{E}(x,-)$ (viewed as sets of morphisms), and the inclusion functor $\mathscr{E}\to{\mathscr{D}}$ induces the following natural isomorphism
$\operatorname{Hom}_{\underline{\mathbb{Z}}_{{\mathscr{D}}}\operatorname{-Mod}}(\mathbb{Z}S,V)\cong\operatorname{Hom}_{\underline{\mathbb{Z}}_{\mathscr{E}}\operatorname{-Mod}}(\mathbb{Z}S,V_{\mathscr{E}})$
for $S\in J_{{\mathscr{D}}}(x)$. Thus formula (6.1) holds if and only if
$\operatorname{Hom}_{\underline{\mathbb{Z}}_{\mathscr{E}}\operatorname{-Mod}}(\mathbb{Z}\mathscr{E}(x,-),V)\cong\operatorname{Hom}_{\underline{\mathbb{Z}}_{\mathscr{E}}\operatorname{-Mod}}(\mathbb{Z}S,V_{\mathscr{E}})$
[…]
Remark 6.10.
Replacing the first condition in the definition of EI categories of type ${\mathbb{N}}$ by the following one: objects in ${\mathscr{C}}$ are parameterized by $n\in\mathbb{Z}$, we can define EI categories of type $\mathbb{Z}$. The reader can mimic our approach to classify Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ where ${\mathscr{C}}$ is an EI category of type $\mathbb{Z}$. Explicitly, we have:
  1. (1)
    Generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by functions
    $\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})=\mathbb{Z}\to{\mathbb{N}}\cup\{\infty\}$
    satisfying the following condition: if $d_{n}\neq 0$, then $d_{n+1}=d_{n}-1$.
  2. (2)
    Non-generic Grothendieck topologies on ${\mathscr{C}^{\mathrm{op}}}$ are parameterized by functions
    $\boldsymbol{d}:\operatorname{Ob}({\mathscr{C}})=\mathbb{Z}\to{\mathbb{N}}\cup\{-\infty\}$
    satisfying the following conditions: $d_{n}=-\infty$ for $n\gg 0$; and if $d_{n}\neq 0$, then $d_{n+1}=d_{n}-1$.
  3. (3)
    A Grothendieck topology $J$ on ${\mathscr{C}^{\mathrm{op}}}$ is rigid if and only if the corresponded function $\boldsymbol{d}$ satisfies the following condition: $d_{n}\neq\infty$ for all $n\in\mathbb{Z}$.
One can deduce the above results from Lemma 5.10. Indeed, let $J$ be a Grothendieck topology on ${\mathscr{C}^{\mathrm{op}}}$. Then for each $n\in\mathbb{Z}$, the full subcategory ${\mathscr{C}}_{\geqslant n}$ of objects $m\geqslant n$ is an ideal of ${\mathscr{C}}$ and an EI category of type ${\mathbb{N}}$. These results follows via applying Propositions 6.3, 6.6 and 6.8 to the restriction of $J$ to ${\mathscr{C}}_{\geqslant n}$. When ${\mathscr{C}}$ is the poset $\mathbb{Z}$, they have been proved in [17, Proposition B.31].
7. Applications
In this section we describe some potential applications of our results in representation theory of combinatorial categories and groups.
7.1. Local self-injectivity of infinite full subcategories
Let ${\mathscr{C}}$ be the category of finite sets and injections, or the category of finite dimensional vector spaces over a finite field $\mathbb{F}_{q}$ and $\mathbb{F}_{q}$-linear injections. It has been proved in [11] (see also [22]) that every indecomposable projective ${\mathscr{C}}$-module over a field $k$ of characteristic 0 is also injective. In this subsection we extend this result to their infinite full subcategories.
Recall that an EI category ${\mathscr{C}}$ of type $\mathbb{N}$ is called locally finite if ${\mathscr{C}}(m,n)$ is a finite set for $m,n\in\mathbb{N}$.
Theorem 7.1.
Let ${\mathscr{C}}$ be a locally finite EI category of type $\mathbb{N}$ such that every morphism is monic, ${\mathscr{D}}$ an infinite full subcategory of ${\mathscr{C}}$, and $k$ a field. If every indecomposable projective ${\mathscr{C}}$-module over $k$ is injective, then so is every indecomposable projective ${\mathscr{D}}$-module over $k$.
Proof.
Since ${\mathscr{D}}$ is a full subcategory of ${\mathscr{C}}$, one may impose a rigid Grothendieck topology $J$ on ${\mathscr{C}}^{{\mathrm{op}}}$ such that $J$-irreducible objects in ${\mathscr{C}}$ are precisely objects in ${\mathscr{D}}$. Thus by Remark 5.17, one has
${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J,\,\underline{k})\simeq{\mathscr{D}}\operatorname{-Mod},$
where $\underline{k}$ is the constant structure sheaf on the site $({\mathscr{C}}^{{\mathrm{op}}},J)$, and ${\mathscr{D}}\operatorname{-Mod}$ is the category of ${\mathscr{D}}$-modules over $k$.
Since ${\mathscr{D}}$ is locally finite, to establish the conclusion, it suffices to show that $k{\mathscr{D}}(n,-)$ is injective for each object $n$ in ${\mathscr{D}}$. Note that by [13, C2.2, Theorem 2.2.3], the equivalence from the sheaf category to the module category is given by the restriction functor; and furthermore the restriction of $P(n)=k{\mathscr{C}}(n,-)$ is precisely $k{\mathscr{D}}(n,-)$. Consequently, one only needs to show that $P(n)$ is an injective object in ${\mathrm{Sh}}({\mathscr{C}}^{{\mathrm{op}}},\,J,\,\underline{k})$. By Corollary 4.4, this is equivalent to saying that $P(n)$ is a $J$-torsion free injective ${\mathscr{C}}$-module. But we have assumed that $P(n)$ is an injective ${\mathscr{C}}$-module, so it remains to show that $P(n)$ is $J$-torsion free.
Since ${\mathscr{C}}$ is an infinite category, $J$ must be a generic Grothendieck topology by Proposition 6.6. Thus $J(n)\subseteq J_{a}(n)$ for $n\in\mathbb{N}$, where $J_{a}$ is the atomic Grothendieck topology. Accordingly, one has $\mathcal{T}(J)\subseteq\mathcal{T}(J_{a})$ by Lemma 3.6, so every $J_{a}$-torsion free ${\mathscr{C}}$-module over $k$ is also $J$-torsion free. Clearly, $P(n)$ is $J_{a}$-torsion free as every morphism in ${\mathscr{C}}$ is monic, so it is $J$-torsion free as well. This finishes the proof. ∎
7.2. Representations of groups
In this subsection we consider applications of sheaf theory in group representation theory. A bridge connecting these two areas is the following construction.
Let $G$ be a group, and $\mathcal{H}$ a family of subgroups of $G$. The orbit category ${\mathrm{Orb}}(G,\mathcal{H})$ of $G$ with respect to $\mathcal{H}$ is defined as follows: objects are left cosets $G/H$ with $H\in\mathcal{H}$, and morphisms from $G/H$ to $G/K$ are $G$-equivariant maps. Every such morphism is an epimorphism, and can be represented by an element $g\in G$ such that $gHg^{-1}\subseteq K$. In particular, for every object $G/H$, its endomorphism monoid is a group isomorphic to $N_{G}(H)/H$, so ${\mathrm{Orb}}(G,\mathcal{H})$ is an EI category. For more details, the reader can refer to [18, III.9]. Moreover, it is easy to see that ${\mathrm{Orb}}(G,\mathcal{H})$ is noetherian (resp., artinian) if and only if the poset $\mathcal{H}$ with respect to inclusion is noetherian (resp., artinian).
Representations of $G$ are closely related to sheaves of modules on ${\mathrm{Orb}}(G,\mathcal{H})$ equipped with certain special Grothendieck topologies. Therefore, one may apply sheaf theory to study representations of groups. We give a few examples appearing in the literature to illustrate this approach.
Example 7.2.
Let $G$ be a topological group, and let $\mathcal{H}$ be a fundamental system of open subgroups. Impose the atomic Grothendieck topology $J_{a}$ on ${\mathrm{Orb}}(G,\mathcal{H})$, and let $\underline{k}$ be a constant structure sheaf given by a commutative ring $k$. Artin’s theorem [18, III.9 Theorem 1] asserts that
${\mathrm{Sh}}({\mathrm{Orb}}(G,\,\mathcal{H}),\,J_{a},\underline{k})\simeq kG\operatorname{-Mod}^{\mathrm{dis}},$
[…]
The following example is introduced by Balmer in [3] and further studied by Xiong, Xu, and Zheng in [26, 27].
Example 7.3.
Let $G$ be a finite group, $\mathcal{H}$ the set of all subgroups of $G$, $p$ a prime number and $k$ a field of characteristic $p$. Following [3, Definition 5.8], we call a family of morphisms $\{G/H_{i}\to G/H\}$ a sipp-covering if there exists a certain $H_{i}$ such that the index $[H:H_{i}]$ is prime to $p$. By [3, Theorem 5.11], these sipp-coverings form a basis of the sipp topology $J$ on ${\mathrm{Orb}}(G,\mathcal{H})$.
Since ${\mathrm{Orb}}(G,\mathcal{H})$ is a finite EI category, $J$ shall be rigid. Actually, an object $G/H$ is $J$-irreducible if and only if $H$ is a $p$-group. Indeed, if $H$ is a $p$-group (including the trivial group) and $\{G/H_{i}\to G/H\}$ is a sipp-covering, then every $H_{i}$ is a $p$-group (otherwise, there is no morphism from $G/H_{i}\to G/H$), and $[H:H_{i}]$ is prime to $p$ for a certain $H_{i}$ if and only if $H$ is conjugate to this $H_{i}$. But this means that $G/H$ is isomorphic to the corresponded $G/H_{i}$. Since ${\mathrm{Orb}}(G,\mathcal{H})$ is an EI category, it follows that all morphisms from this $G/H_{i}$ to $G/H$ are isomorphisms, so $G/H$ is a $J$-irreducible object.
On the other hand, suppose that $H$ is not a $p$-group and let $H_{p}$ be its Sylow $p$-subgroup. Then any morphism $G/H_{p}\to G/H$ forms a sipp-covering. It generates a sieve contained in $J(G/H)$ which is not the maximal sieve. Consequently, by letting $\mathscr{D}$ be the full subcategory of ${\mathrm{Orb}}(G,\mathcal{H})$ consisting of objects $G/H$ with $H$ a $p$-subgroup, one obtains the following equivalence
${\mathrm{Sh}}({\mathrm{Orb}}(G,\mathcal{H}),\,J,\,\underline{k})\simeq k{\mathscr{D}}^{{\mathrm{op}}}\operatorname{-Mod},$
a result described in [26, Subsection 5.3].
A further generalization of the above example is:
Example 7.4.
Let $G$ be an artinian group (for instance, the Prüfer $p$-group), $\mathcal{H}$ a set of subgroups of $G$, and $k$ a commutative ring. The orbit category ${\mathrm{Orb}}(G,\mathcal{H})$ is an artinian EI category, so by Theorem 5.7 every Grothendieck topology $J$ on it is rigid. By Theorem 5.16, one has
${\mathrm{Sh}}({\mathrm{Orb}}(G,\mathcal{H}),\,J,\,\underline{k})\simeq k{\mathscr{D}}^{{\mathrm{op}}}\operatorname{-Mod},$
where ${\mathscr{D}}$ is the full subcategory of ${\mathrm{Orb}}(G,\mathcal{H})$ consisting of $J$-irreducible objects.
There already exist quite a few reformulations of Alperin’s weight conjecture. In particular, Webb gives in [24] a reformulation in terms of the standard stratification property of the category algebra of the orbit category. The previous example gives a sheaf theoretic reformulation. We end this paper with an illustration of this observation.
Let $G$ be a finite group, $p$ a prime number, and $k$ an algebraically closed field of characteristic $p$. A weight is a pair $(H,L)$ with $H$ a $p$-subgroup (possibly trivial) of $G$ and $L$ is an irreducible projective $kN_{G}(H)/H$-module. Two weights $(H_{1},L_{1})$ and $(H_{2},L_{2})$ are equivalent if there exists $g\in G$ such that the conjugate $H_{1}^{g}$ coincides with $H_{2}$ and $L_{1}^{g}\cong L_{2}$. Alperin’s weight conjecture asserts that the number of isomorphism classes of irreducible $kG$-modules is equal to the number of equivalence classes of weights.
Let $\mathcal{H}$ be the family of $p$-subgroups of $G$ and equip ${\mathrm{Orb}}(G,\mathcal{H})$ with the atomic Grothendieck topology $J_{a}$. Since only the object $G/1$ is $J_{a}$-irreducible, and $N_{G}(1)/1\cong G\cong G^{{\mathrm{op}}}$, one obtains
${\mathrm{Sh}}({\mathrm{Orb}}(G,\mathcal{H}),\,J,\,\underline{k})\simeq kG\operatorname{-Mod}.$
Therefore, Alperin’s weight conjecture can be reformulated as follows: the number of isomorphism classes of irreducible sheaves is equal to the number of equivalence classes of weights.
Since the automorphism group of $G/H$ is $N_{G}(H)/H$, a weight of $G$ can be viewed as a “local” representation of ${\mathrm{Orb}}(G,\mathcal{H})^{{\mathrm{op}}}$; that is, a representation supported only on objects isomorphic to $G/H$. Instead, irreducible objects in ${\mathrm{Sh}}({\mathrm{Orb}}(G,\mathcal{H}),\,J,\,\underline{k})$ can be viewed as “global” representations of ${\mathrm{Orb}}(G,\mathcal{H})^{{\mathrm{op}}}$. Thus this reformulation reflects the “local-global” principle of Alperin’s weight conjecture. Furthermore, one may form a stronger version of this conjecture as follows: construct a bidirectional machinery such that from each weight one can obtain a unique irreducible sheaf up to isomorphism, and by a converse procedure one recovers this weight from the isomorphism class of the irreducible sheaf.
Let ${\mathscr{C}}={\mathrm{Orb}}(G,\mathcal{H})^{{\mathrm{op}}}$ and $k{\mathscr{C}}$ the category algebra of ${\mathscr{C}}$. Note that “local” representations and “global” representations of ${\mathscr{C}}$ are related by the restriction and induction functors. Thus one may ask the following questions:
  1. (1)
    For each weight $(H,L)$, can we find a unique $J$-saturated irreducible ${\mathscr{C}}$-module $V(H,L)$ which is a composition factor of
    $k{\mathscr{C}}\otimes_{k(N_{G}(H)/H)}L$
    such that $V(H,L)\cong V(H^{\prime},L^{\prime})$ if and only if $(H,L)$ and $(H^{\prime},L^{\prime})$ are equivalent?
  2. (2)
    If the answer of the above question is yes, can every irreducible $J_{a}$-saturated ${\mathscr{C}}$-module be obtained via the above construction?
If one can answer these two questions affirmatively, then we can get a bijective correspondence between the number of equivalence classes of weights and the set of isomorphism classes of irreducible $J_{a}$-saturated $k{\mathscr{C}}$-modules. A proof of Alperin’s wight conjecture then follows from this bijection.